### v1

# Data-based control of Logical Networks

Abstract
In recent years, data-driven approaches have become increasingly pervasive across all areas of control engineering. However, the applications of data-based techniques to Boolean Control Networks (BCNs) are still very limited. In this paper we aim to fill this gap, by exploring the possibility of solving three fundamental control problems, i.e., state feedback stabilization, safe control and output regulation, for a BCN, leveraging only a limited amount of data generated by the network, without knowing or identifying its model.
keywords:
Boolean control networks; feedback stabilization; safe control; output regulation.
thanks: Corresponding Author.
1 Introduction
Boolean Networks (BNs) are discrete-time (autonomous) dynamical systems whose describing variables are constrained to take values in the Boolean set. When their evolution is controlled by an external signal (an input), whose entries in turn take Boolean values, they are called Boolean Control Networks (BCNs). The interest in this class of logical systems originated in the 1960s, when Stuart Kauffman introduced them to model and analyze gene regulatory networks Kauffman (1969), and even if this may still be regarded as their most successful application area, they have been adopted in a wide range of fields, including systems biology, computer science, game theory, smart cities and network theory (see, e.g., Bornholdt (2008); Cheng (2014); Green et al. (2007); Kabir et al. (2014)).
The algebraic approach to BNs and BCNs, proposed by Cheng et al. Cheng (2009); Cheng & Qi (2010a, b); Cheng et al. (2011), and deeply relying on the concept of semi-tensor product (STP) Cheng (2001), has made it possible to represent BNs and BCNs by means of discrete-time state space models in which the state, input and output are logical vectors, and the describing matrices are logical matrices, in turn. In this way, concepts and tools developed for linear time-invariant state space models have been adjusted to formalize and solve control problems for BNs and BCNs (see, e.g., Fornasini & Valcher (2014); Li et al. (2013)).
Despite the simplifying “on/off" (“active/inactive", “high/low") logic at the basis of BCNs, the size and complexity of the physical systems they model (e.g., gene regulatory networks) make it difficult to derive an accurate model of the network, as well as to obtain it by means of identification techniques. On the other hand, in recent times direct data-driven methods have gained interest in the control community, given their attractive feature of bypassing the identification phase, to directly design the problem solution based on raw data. Yet, the use of partial data to solve control problems for logical networks is still at an early stage, and to the best of our knowledge the first and only attempt in this direction has been made in Leifeld et al. (2018), where a one-step data-driven approach to solve the output stabilization problem is proposed. More specifically, the authors in Leifeld et al. (2018) first reconstruct from the available data the output prediction matrix that maps the current input/output pair to the set of all possible next outputs, and then, based on this map, design an output feedback law to stabilize the output of the BCN to a desired value. However, in general, the set of all possible next outputs contains more than one element and hence the output prediction is possibly non deterministic. This uncertainty is handled by resorting to Probabilistic Boolean Networks (PBNs) to compensate for the fact that data are not sufficient to identify the whole BCN and uniquely solve the stabilization problem.
In this paper, we adopt a different viewpoint and exploit the concept of informativity introduced by van Waarde, Eising, and coauthors in van Waarde et al. (2020) (see, also, Waarde et al. (2023)), to address the limitations arising from having only a limited amount of data available, without the need to resort to PBNs. Indeed, the informativity framework is based on the idea that, when data are not informative enough to uniquely identify the system, the best that we can do is imposing that all the systems compatible with the collected data (among which there is surely also our target system) possess the desired property or behave in a certain way. Therefore, in this work we solve three fundamental control problems, namely stabilization, safe control and output regulation of a BCN by state feedback, by solving them for all the logical systems compatible with the available data. This automatically guarantees that the problem is solved for the actual BCN, even under partial information. More specifically, we provide necessary and sufficient conditions on the collected data to stabilize a BCN to a certain state, to drive its evolution towards a safe set, and to set its output to a desired value, by means of a state feedback. Moreover, for each of these problems, we propose a procedure to design such a feedback matrix based on the collected data.
The paper is organized as follows. Section 2 contains some preliminary notions and results about BNs and BCNs. In Section 3 we formalize and solve the data-driven feedback stabilization problem for a BCN. The safe control problem is investigated in Section 4. In Section 5 we state and solve the data-driven output regulation problem by state feedback for a BCN. Section 6 concludes the paper.
This paper extends a manuscript recently submitted to CDC 2025. Compared to the conference version the main novel contribution is represented by Section 4.
Notation. Given two nonnegative integers $k$ and $n$, with $k\leq n$, we denote by $[k,n]$ the set of integers $\{k,k+1,\dots,n\}$. We consider Boolean vectors and matrices whose entries take values in $\mathcal{B}\triangleq\{0,1\}$, equipped with the usual logical operations, namely sum (OR) $\vee$, product (AND) $\land$ and negation $\neg$. We denote by $\delta_{k}^{i}$ the $k$-dimensional $i$-th canonical vector, by $\mathcal{L}_{k}$ the set of all $k$-dimensional canonical vectors, and by $\mathcal{L}_{k\times q}\subset\mathcal{B}^{k\times q}$ the set of all $k\times q$ logical matrices, namely matrices whose $q$ columns are canonical vectors of size $k$. Any matrix $L\in\mathcal{L}_{k\times q}$ can be described as $L=[\delta_{k}^{i_{1}}\ \delta_{k}^{i_{2}}\ \dots\ \delta_{k}^{i_{q}}]$, for some indices $i_{1},i_{2},\dots,i_{q}\in[1,k]$. $I_{n}$ denotes the $n$-dimensional identity matrix. The $k$-dimensional vector with all zero entries is denoted by $\mathbbold{0}_{k}$, and the $k\times l$ matrix with all zero entries by $\mathbbold{0}_{k\times l}$. The $(i,j)$-th entry of a matrix $M$ is denoted by $[M]_{ij}$, and the $j$-th entry of a vector ${\bf v}$ by $[{\bf v}]_{j}$. The $j$-th column of matrix $M$ is denoted by ${\rm col}_{j}(M)$. We denote by $\odot$ the Hadamard (or entry-wise) product. A Boolean (in particular, a logical) square matrix $L$ of size $k$ is reducible if there exists a permutation matrix $P\in\mathcal{L}_{k\times k}$ such that
$P^{\top}LP=\begin{bmatrix}L_{11}&L_{12}\cr 0&L_{22}\end{bmatrix},$
where $L_{11}$ and $L_{22}$ are square Boolean (logical) matrices. Otherwise, $L$ is called irreducible. $L$ is irreducible if and only if $I_{k}\vee L\vee\dots\vee L^{k-1}$ has all unitary entries.
There is a bijective correspondence between Boolean variables $X\in\mathcal{B}$ and vectors ${\bf x}\in\mathcal{L}_{2}$, defined by ${\bf x}=[X\ \neg X]^{\top}$. Given $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{p\times q}$, their (left) semi-tensor product ($\ltimes$) (see Cheng (2001)) is
$A\ltimes B\triangleq(A\otimes I_{l/n})(B\otimes I_{l/p}),\quad l\triangleq{\rm l.c.m.}\{n,p\}.$
For the properties of the semi-tensor product we refer to Cheng et al. (2011). By exploiting the semi-tensor product, it is possible to extend the previously mentioned bijective correspondence to a bijective correspondence between $\mathcal{B}^{n}$ and $\mathcal{L}_{2^{n}}$, as follows. Given a vector $X=[X_{1}\ X_{2}\ \dots\ X_{n}]^{\top}\in\mathcal{B}^{n}$, set
${\bf x}\triangleq\begin{bmatrix}X_{1}\\ \neg X_{1}\end{bmatrix}\ltimes\begin{bmatrix}X_{2}\\ \neg X_{2}\end{bmatrix}\ltimes\dots\ltimes\begin{bmatrix}X_{n}\\ \neg X_{n}\end{bmatrix}.$
It is possible to associate with a Boolean (in particular, a logical) matrix $L\in{\mathcal{B}}^{k\times k}$ a directed graph (or, digraph), $\mathcal{D}(L)=(\mathcal{V},\mathcal{E})$, where ${\mathcal{V}}=[1,k]$ is the set of nodes and ${\mathcal{E}}$ is the set of edges. An edge $(j,l)$, from $j$ to $l$, belongs to $\mathcal{E}$ if and only if $[L]_{lj}=1$. A sequence $j_{1}\to j_{2}\to\dots\to j_{r}\to j_{r+1}$ in $\mathcal{D}(L)$ is a path of length $r$ from $j_{1}$ to $j_{r+1}$ if $(j_{1},j_{2}),\dots,(j_{r},j_{r+1})$ are edges in $\mathcal{E}$. A closed path is a cycle. A node $j^{*}\in{\mathcal{V}}$ is said to be globally reachable if there exists a path to $j^{*}$ from any other node in the network. A digraph $\mathcal{D}(L)$ is strongly connected if all its nodes are globally reachable, and this is the case if and only if $L$ is irreducible.
2 Boolean Control Networks: Preliminaries
A Boolean control network (BCN) is a logical system described by the following equations
$X(t+1)$ $=$ $f(X(t),U(t)),\quad t\in\mathbb{Z}_{+},$ (1a)
$Y(t)$ $=$ $h(X(t)),$ (1b)
where $X(t)\in\mathcal{B}^{n}$ is the $n$-dimensional state variable, $U(t)\in\mathcal{B}^{m}$ is the $m$-dimensional input, and $Y(t)\in\mathcal{B}^{p}$ is the $p$-dimensional output at time $t$. $f$ and $h$ are logical functions, i.e., $f:\mathcal{B}^{n}\times\mathcal{B}^{m}\to\mathcal{B}^{n}$, while $h:\mathcal{B}^{n}\to\mathcal{B}^{p}$. If the logical system is autonomous, namely no input acts on it, then the BCN becomes a Boolean network (BN) and equation (1a) becomes
$X(t+1)=f(X(t)),\quad t\in\mathbb{Z}_{+},$ (2)
where $f:\mathcal{B}^{n}\to\mathcal{B}^{n}$ is a logical function. By exploiting the bijective correspondence between Boolean and logical vectors, the BN (2) can be equivalently expressed via its algebraic representation Cheng et al. (2011) as
${\bf x}(t+1)=L{\bf x}(t),\quad t\in\mathbb{Z}_{+},$ (3)
where ${\bf x}(t)\in\mathcal{L}_{N}$ and $L\in{\mathcal{L}}_{N\times N}$, with $N=2^{n}$. Similarly, the algebraic representation of the BCN (1) is
${\bf x}(t+1)$ $=$ $L\ltimes{\bf u}(t)\ltimes{\bf x}(t),\quad t\in\mathbb{Z}_{+},$ (4a)
${\bf y}(t)$ $=$ $H{\bf x}(t),$ (4b)
where ${\bf x}(t)\in\mathcal{L}_{N}$, ${\bf u}(t)\in\mathcal{L}_{M}$, ${\bf y}(t)\in\mathcal{L}_{P}$, $L\in\mathcal{L}_{N\times NM}$ and $H\in{\mathcal{L}}_{P\times N}$, with $N=2^{n}$, $M=2^{m}$ and $P=2^{p}$. It is worth remarking that all the subsequent analysis holds for generic nonnegative integer values of $N,M$ and $P$, not necessarily powers of $2$. The matrix $L\in\mathcal{L}_{N\times NM}$, whose columns are canonical vectors of size $N$, can be divided into $M$ square blocks of size $N$ as follows
$L=\left[\begin{array}[]{c|c|c|c}L_{1}&L_{2}&\dots&L_{M}\end{array}\right],$ (5)
where each block $L_{i}\in\mathcal{L}_{N\times N},i\in[1,M]$, describes the behavior of the $i$-th subsystem of the BCN, i.e., the Boolean network obtained when ${\bf u}(t)=\delta_{M}^{i},\forall t\in\mathbb{Z}_{+}$:
${\bf x}(t+1)=L_{i}{\bf x}(t),\quad t\in\mathbb{Z}_{+}.$ (6)
Definition 1.
Cheng et al. (2011) Given a BCN as in (4a), we say that ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if there exist $\tau\in\mathbb{Z}_{+}$ and an input ${\bf u}(t),t\in[0,\tau-1],$ that leads the state trajectory from ${\bf x}(0)={\bf x}_{0}$ to ${\bf x}(\tau)={\bf x}_{f}$. The whole BCN is reachable if the previous condition holds for every ${\bf x}_{0}$ and ${\bf x}_{f}$ in ${\mathcal{L}}_{N}$.
We can characterize the reachability of a single state and of the whole network Cheng et al. (2011) by resorting to the matrix
$L_{tot}\triangleq L_{1}\vee L_{2}\vee\dots\vee L_{M}\in{\mathcal{B}}^{N\times N},$ (7)
and the associated digraph $\mathcal{D}(L_{tot})\triangleq(\mathcal{V},\mathcal{E}_{tot})$. The state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if and only if there exists $\tau\in\mathbb{Z}_{+}$ such that $[L_{tot}^{\tau}]_{ij}=1$, or equivalently, there is a path of length $\tau$ from $j$ to $i$ in $\mathcal{D}(L_{tot})$. Consequently, a BCN is reachable if and only if the matrix $L_{tot}$ is irreducible, or equivalently $\mathcal{D}(L_{tot})$ is strongly connected. In the following, we will denote by $\mathcal{R}({\bf x}_{0})$ the set of states reachable from ${\bf x}_{0}$.
Definition 2.
Given a BN as in (3), a state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ is an equilibrium point of the BN if ${\bf x}_{e}=L{\bf x}_{e}$. Given a BCN as in (4a), we say that ${\bf x}_{e}$ is an equilibrium point of the BCN corresponding to the input ${\bf u}=\delta^{i}_{M}$ if ${\bf x}_{e}$ is an equilibrium point of its $i$-th subsystem (6), i.e., ${\bf x}_{e}=L\ltimes\delta^{i}_{M}\ltimes{\bf x}_{e}$.
Note that ${\bf x}_{e}=\delta^{j}_{N}$ is an equilibrium point of a BN or a BCN if and only if the node $j$ has a self-loop in ${\mathcal{D}}(L)$ or ${\mathcal{D}}(L_{tot})$, respectively.
Definition 3.
Cheng & Liu (2009); Cheng et al. (2011); Fornasini & Valcher (2013b) A BCN (4a) is stabilizable to the state ${\bf x}_{e}\in\mathcal{L}_{N}$ if for every ${\bf x}(0)\in\mathcal{L}_{N}$ there exist $\tau\in\mathbb{Z}_{+}$ and ${\bf u}(t),t\in\mathbb{Z}_{+}$, that leads to ${\bf x}(t)={\bf x}_{e},\forall t\geq\tau$.
Stabilizability can be characterized as follows.
Proposition 4.
Cheng et al. (2011); Fornasini & Valcher (2013b); Li et al. (2013) A BCN (4a) is stabilizable to ${\bf x}_{e}\in\mathcal{L}_{N}$ if and only if the following conditions hold:
  • (i)
    ${\bf x}_{e}$ is an equilibrium point of the BCN corresponding to some input ${\bf u}=\delta^{i}_{M}$, $i\in[1,M]$;
  • (ii)
    ${\bf x}_{e}$ is reachable from every state ${\bf x}(0)$, i.e., ${\bf x}_{e}\in\bigcap_{{\bf x}(0)\in\mathcal{L}_{N}}\mathcal{R}({\bf x}(0))$ (i.e., ${\bf x}_{e}$ is globally reachable).
If we define the basin of attraction of a state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ as the set of states ${\bf x}\in{\mathcal{L}}_{N}$ from which it is possible to reach ${\bf x}_{e}$ and denote it by ${\mathcal{S}}({\bf x}_{e})$, then condition (ii) in Proposition 4 is equivalent to ${\mathcal{S}}({\bf x}_{e})={\mathcal{L}}_{N}$.
By Proposition 4, in order to identify the states to which system (4a) can be stabilized, we first need to determine the set $\mathcal{X}_{e}$ of the equilibrium points of the BCN (corresponding to the various inputs), namely the nodes of $\mathcal{D}(L_{tot})$ with a self-loop. Then, we need to identify the states in $\mathcal{X}_{e}$ that are reachable from every other state, namely the states in $\mathcal{X}_{e}$ whose associated node in $\mathcal{D}(L_{tot})$ is globally reachable. We will henceforth interchangeably write that a node (in the digraph) or a state (of the BCN) are globally reachable.
It has been shown in Cheng et al. (2011); Fornasini & Valcher (2013b); Li et al. (2013) that if a BCN is stabilizable to some ${\bf x}_{e}$, then it is stabilizable by means of a state feedback law: ${\bf u}(t)=K{\bf x}(t),\exists K\in{\mathcal{L}}_{M\times N}.$ When so the BCN becomes a BN:
${\bf x}(t+1)=L\ltimes K\ltimes{\bf x}(t)\ltimes{\bf x}(t)=\tilde{L}(K){\bf x}(t),$ (8)
for some $\tilde{L}(K)\in{\mathcal{L}}_{N\times N}.$
3 Data-driven stabilization
In this section we restrict our attention to the state equation (4a) of the BCN. We assume to have performed an offline experiment during which we have collected some state/input data from the BCN (4a) on a finite time interval $[0,T],T\in\mathbb{Z}_{+}$111 Instead of a single experiment to collect data, we can perform several experiments, say $r$, on some finite time intervals $[0,T_{i}],i\in[1,r]$. It suffices to arrange data consistently with the way it is illustrated in the case of a single experiment., with $T$ sufficiently large (in a sense that will be clarified in the sequel), and we define the vector sequences ${\bf x}_{d}\triangleq\{{\bf x}_{d}(t)\}_{t=0}^{T}$, and ${\bf u}_{d}\triangleq\{{\bf u}_{d}(t)\}_{t=0}^{T-1}$. We rearrange the collected data into the following matrices:
$X_{p}$$\triangleq$ $\begin{bmatrix}{\bf x}_{d}(0)&{\bf x}_{d}(1)&\dots&{\bf x}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{N\times T},$ (9)
$X_{f}$$\triangleq$ $\begin{bmatrix}{\bf x}_{d}(1)&{\bf x}_{d}(2)&\dots&{\bf x}_{d}(T)\end{bmatrix}\in\mathcal{L}_{N\times T},$ (10)
$U_{p}$$\triangleq$ $\begin{bmatrix}{\bf u}_{d}(0)&{\bf u}_{d}(1)&\dots&{\bf u}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{M\times T},$ (11)
where the subscripts $p$ and $f$ stand for past and future, respectively.
As a BCN (4a) bijectively corresponds to its describing matrix, we define the set of Boolean control networks (4a) compatible with the data $({\bf x}_{d}$, ${\bf u}_{d})$ as
${\mathcal{B}}_{d}$$\triangleq$ $\{\tilde{L}\in\mathcal{L}_{N\times NM}:$
${\bf x}_{d}(t+1)=\tilde{L}\ltimes{\bf u}_{d}(t)\ltimes{\bf x}_{d}(t),\forall t\in[0,T-1]\}.$
Obviously, the matrix $L$ of the BCN that has generated the data belongs to ${\mathcal{B}}_{d}$. However, ${\mathcal{B}}_{d}$ may also contain other matrices if the data do not allow to uniquely determine $L$. To this end, we introduce the notion of informativity for identifiability.
Definition 5.
We say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of the BCN (4a) if ${\mathcal{B}}_{d}\equiv\{L\}$.
We can provide the following characterization of informativity for identifiability.
Lemma 6.
The data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of the BCN (4a) if and only if the logical matrix
$\left[\begin{array}[]{c|c|c}{\rm col}_{1}(U_{p})\ltimes{\rm col}_{1}(X_{p})&\dots&{\rm col}_{T}(U_{p})\ltimes{\rm col}_{T}(X_{p})\end{array}\right]$ (13)
is of full row rank.
Proof.
A BCN (4a) is uniquely identified from the data matrices $X_{p}$, $X_{f}$ and $U_{p}$ if and only if these latter allow to deduce each column of the matrix $L$ in (4a), or, equivalently, are representative of every possible transition that can occur in the network (from all possible state vectors and corresponding to all possible inputs). By taking into account that the $k$-th column of the matrix $X_{f}$ represents the successor state of the state given by the $k$-th column of $X_{p}$, when the applied input is the one corresponding to the $k$-th column of $U_{p}$, it follows that we know from data every possible transition of the BCN if and only if for every $(i,j)\in[1,M]\times[1,N]$ there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr U_{p}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{M}^{i}\end{bmatrix},$ (14)
which, by resorting to the semi-tensor product, is equivalent to ${\rm col}_{k}(U_{p})\ltimes{\rm col}_{k}(X_{p})=\delta_{M}^{i}\ltimes\delta_{N}^{j}.$ But, saying that for every $(i,j)\in[1,M]\times[1,N]$ condition (14) holds for some $k\in[1,T]$ amounts to saying that the matrix (13) contains all vectors of $\mathcal{L}_{NM}$, i.e., is of full row rank. ∎
Remark 7.
If (14) holds, then $X_{f}\delta_{T}^{k}={\rm col}_{j}(L_{i}).$
To make the following analysis meaningful, from now on we assume that the collected data are not informative for identifiability of the considered BCN, namely ${\mathcal{B}}_{d}\supsetneq\{L\}$, and we investigate under what conditions it is still possible to use them to solve some control problems.
We now formalize the data-driven state feedback stabilization problem.
Problem 1.
Given the data $({\bf x}_{d}$, ${\bf u}_{d})$, determine (if it exists) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that ${\bf u}(t)=K{\bf x}(t)$ stabilizes all the BCNs compatible with the data to some state ${\bf x}_{e}\in\mathcal{L}_{N}$.
When Problem 1 is solvable, we will say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for stabilization to ${\bf x}_{e}\in\mathcal{L}_{N}$ by state feedback.
By Proposition 4, it is possible to stabilize a BCN to some state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ only if such a state is an equilibrium point of the BCN. Therefore Problem 1 is solvable only if ${\bf x}_{e}$ is an equilibrium point for all possible BCNs compatible with the data, namely all the BCNs in ${\mathcal{B}}_{d}$. By Definition 2, a state ${\bf x}_{e}$ is an equilibrium point of a BCN (4a) if there exists $i\in[1,M]$ such that for ${\bf u}=\delta_{M}^{i}$ the successor of ${\bf x}_{e}$ is ${\bf x}_{e}$ itself. Therefore, in order to identify the equilibrium points compatible with the collected data, we can stack the matrices $X_{p}$ and $X_{f}$ one above the other, and check in which columns the logical vector in the upper matrix is the same as the logical vector in the lower matrix. Such a logical vector is an equilibrium point, since the column of $X_{f}$ represents the successor of the corresponding column in $X_{p}$ (for some input). More formally, we can apply the following algorithm.
Algorithm 1: Equilibrium points of the BCNs in ${\mathcal{B}}_{d}$
Input: The data matrices $X_{p}$ and $X_{f}$.
Output: The set $\mathcal{X}_{e}^{d}$ of all equilibrium points ${\bf x}_{e}$ compatible with the data.
Initialization:
Set $\bar{X}=\mathbbold{0}_{N\times T}$.
Procedure:
for $(j,k)\in[1,N]\times[1,T]$, do
  • $[\bar{X}]_{jk}=[X_{p}]_{j,k}[X_{f}]_{j,k}$
end for
Set:
  • $\{(j_{1},k_{1}),(j_{2},k_{2}),\dots,(j_{n_{e}},k_{n_{e}})\}$ = nonzero$(\bar{X})$
  • $\mathcal{X}_{e}^{d}=\{\delta_{N}^{j_{1}},\delta_{N}^{j_{2}},\dots,\delta_{N}^{j_{n_{e}}}\}$.
Conclusion: Return $\mathcal{X}_{e}^{d}$. ∎
Note that Algorithm 1 first calculates the Hadamard product $\bar{X}=X_{p}\odot X_{f}$ of the logical matrices $X_{p}$ and $X_{f}$, and then identifies the $n_{e}\geq 0$ positions of the nonzero entries of $\bar{X}$: the first (i.e., row) indices of these positions are the indices of the logical vectors that represent equilibrium points of all the BCNs in ${\mathcal{B}}_{d}$.
Once we have determined the equilibria in $\mathcal{X}_{e}^{d}$, we can identify the input(s) to which they correspond. Indeed, if ${\bf x}_{e}=\delta_{N}^{j}\in\mathcal{X}_{e}^{d}$, it means that $\exists k\in[1,T]$ such that
$\bar{X}\delta_{T}^{k}=\delta_{N}^{j}\quad\Leftrightarrow\quad\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{j}\end{bmatrix}.$
Therefore, ${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}=\delta_{M}^{i},\exists i\in[1,M]$, is the/a corresponding input. Once we know $\mathcal{X}_{e}^{d}$, we can verify if the state ${\bf x}_{e}$ to which we want to stabilize the BCNs in ${\mathcal{B}}_{d}$ belongs to $\mathcal{X}_{e}^{d}$. If not, the problem is not solvable, while if ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$ we need to verify (using only data) that condition (ii) of Proposition 4 holds, namely that ${\bf x}_{e}$ is globally reachable, in all the networks in ${\mathcal{B}}_{d}$. To this end, we introduce the notion of informativity for reachability of a generic $\bar{\bf x}\in{\mathcal{L}}_{N}$.
Definition 8.
Given $\bar{\bf x}\in\mathcal{L}_{N}$, we say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of $\bar{\bf x}$ if $\bar{\bf x}$ is globally reachable in all the BCNs in ${\mathcal{B}}_{d}$.
To test informativity for reachability, we propose the following algorithm.
Algorithm 2: Reachability of $\bar{\bf x}\in\mathcal{L}_{N}$
Input: - The data matrices $X_{p},X_{f},$ and $U_{p}$;
           - $\bar{\bf x}=\delta_{N}^{j}\in\mathcal{L}_{N}$.
Output: - Is $\bar{\bf x}$ globally reachable for every BCN in ${\mathcal{B}}_{d}$?
               Yes/No.
             - The basin of attraction $\mathcal{S}(\bar{\bf x})$ of $\bar{\bf x}$.
Initialization:
Set $d=0$, ${\mathcal{S}}_{0}=\{\delta_{N}^{j}\}$, $({\bf u}_{1}^{\bar{\bf x}},{\bf u}_{2}^{\bar{\bf x}},\dots,{\bf u}_{N}^{\bar{\bf x}})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
Iterative procedure:
if ${\mathcal{S}}_{d}\neq\emptyset$, then
  • $d\leftarrow d+1$
  • ${\mathcal{S}}_{d}=\emptyset$
  • for $k\in[1,T]$, do
            if
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix},\qquad\exists\ \delta^{q}_{N}\in{\mathcal{S}}_{d-1}$
    and $\delta^{r}_{N}\not\in\cup_{i=1}^{d}{\mathcal{S}}_{i}$, then
                   ${\mathcal{S}}_{d}\leftarrow{\mathcal{S}}_{d}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}^{\bar{\bf x}}\triangleq U_{p}\delta_{T}^{k}$
            end if
    end for
  • Go back to the Iterative procedure.
else set $\mathcal{S}(\bar{\bf x})=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$ and go to the Conclusion
end if
Conclusion: if $\mathcal{S}(\bar{\bf x})=\mathcal{L}_{N}$, then return (‘Yes’, $\mathcal{S}(\bar{\bf x})$)
                 else return (‘No’, $\mathcal{S}(\bar{\bf x})$)
                 end if
Note that the Algorithm identifies first the set of states (different from $\bar{\bf x}=\delta_{N}^{j}$) that have $\bar{\bf x}=\delta_{N}^{j}$ as their successor, namely ${\mathcal{S}}_{1}$. For each such state $\delta^{r}_{N}$, it memorizes in ${\bf u}_{r}^{\bar{\bf x}}$ one of the values of the input that allows the transition from $\delta^{r}_{N}$ to $\bar{\bf x}=\delta^{j}_{N}$. Then it identifies the set ${\mathcal{S}}_{2}$ of the states that do not belong to ${\mathcal{S}}_{0}\cup{\mathcal{S}}_{1}=\{\bar{\bf x}\}\cup{\mathcal{S}}_{1}$, but have a successor in ${\mathcal{S}}_{1}$. Again, for each $\delta^{r}_{N}\in{\mathcal{S}}_{2}$, it memorizes one of the values of the input that allows the transition from $\delta^{r}_{N}$ to some state in ${\mathcal{S}}_{1}$. Since the set of states of a BCN is finite, there exists $d\geq 1$ such that ${\mathcal{S}}_{d}=\emptyset$. If at that stage $\mathcal{S}(\bar{\bf x})=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}=\mathcal{L}_{N}$, this means that $\bar{\bf x}=\delta^{j}_{N}$ is globally reachable and $\mathcal{S}(\bar{\bf x})$ is its basin of attraction.
We are now ready to state Lemma 9, whose proof is straightforward from the previous analysis.
Lemma 9.
Given $\bar{\bf x}\in\mathcal{L}_{N}$, the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of $\bar{\bf x}$ if and only if Algorithm 2 terminates with an affermative answer.
By putting together the results obtained so far, and leveraging the model-based results in Proposition 4 and subsequent comments, we can provide the solution to Problem 1, namely a characterization of informativity for stabilization by state feedback.
Theorem 10.
Problem 1 is solvable (i.e., the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for stabilization of ${\bf x}_{e}$ by state feedback) if and only if
  • (i)
    ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$;
  • (ii)
    the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of ${\bf x}_{e}$.
Moreover, the $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by  ${\rm col}_{r}(K)={\bf u}_{r}^{{\bf x}_{e}},\forall r\in[1,N],$ where ${\bf u}_{r}^{{\bf x}_{e}}$ is the logical vector determined in Algorithm 2.
Proof.
(Necessity). Assume, by contradiction, that condition (i) does not hold, namely ${\bf x}_{e}\notin\mathcal{X}_{e}^{d}$. This means that there exists at least one BCN in $\mathcal{B}_{d}$ for which ${\bf x}_{e}$ is not an equilibrium point and hence, by Proposition 4, such network cannot be stabilized to ${\bf x}_{e}$. This contradicts the fact that Problem 1 is solvable. Assume now that ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$, but condition (ii) fails to hold. This means that there exists at least one BCN in $\mathcal{B}_{d}$ for which ${\bf x}_{e}$ is not globally reachable. One of such BCNs is simply obtained by imposing that all the transitions that cannot be deduced from data are self-loops. By Proposition 4, such BCNs cannot be stabilized, and hence again Problem 1 is not solvable.
(Sufficiency). It follows directly from Proposition 4 and previous comments and results.
The final statement follows from the fact that, for every state $\delta_{N}^{r}$ of every BCN in ${\mathcal{B}}_{d}$, ${\bf u}_{r}^{{\bf x}_{e}}$ is exactly the/an input that, starting from this state, ensures the correct transition towards reaching ${\bf x}_{e}$, by construction (see comments after Algorithm 2). ∎
Remark 11.
A necessary condition for the data $({\bf x}_{d}$, ${\bf u}_{d})$ to be informative for stabilization to ${\bf x}_{e}\in\mathcal{L}_{N}$ by state feedback is that the matrix $X_{p}$ is of full row rank [^2] This imposes $T\geq N$., or, equivalently, each state of $\mathcal{L}_{N}$ appears at least once as a column of $X_{p}$. Indeed, in order to check that the state ${\bf x}_{e}=\delta^{j}_{N}$ belongs to $\mathcal{X}_{e}^{d}$, namely is an equilibrium point of all BCNs compatible with the data, there must be a column of $X_{p}$ that coincides with ${\bf x}_{e}$ (and whose corresponding column in $X_{f}$ also coincides with ${\bf x}_{e}$). On the other hand, if there is some state, say $\delta_{N}^{q}$, $q\neq j$, that does not appear as a column of $X_{p}$, we cannot deduce from the data its successor for any $i$-th input, $i\in[1,M]$. This implies that in all the BCNs compatible with the data, the successors of $\delta_{N}^{q}$ can be arbitrarily chosen. As a consequence, in all the BCNs in $\mathcal{B}_{d}$ for which ${\rm col}_{q}(\tilde{L}_{i})=\delta_{N}^{q},\forall i\in[1,M]$, ${\bf x}_{e}$ would not be globally reachable. This prevents the data $({\bf x}_{d}$, ${\bf u}_{d})$ to be informative for reachability of ${\bf x}_{e}$.
4 Data-driven safe control
In Fornasini & Valcher (2016) the safe control of a BCN was first addressed, then it was extended to handle random impulsive logical control networks in Zhou et al. (2023) and probabilistic Boolean control networks using an event-triggered approach in Shao et al. (2025). The safe control problem is stated as follows. Given a BCN (4a), suppose that the set of states ${\mathcal{L}}_{N}$ can be partitioned into a set of unsafe states ${\mathcal{X}}_{u}$ and a set of safe states ${\mathcal{X}}_{s}\triangleq{\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}$. Under what conditions is it possible to design a control input so that every state trajectory that originates from a safe state remains indefinitely in ${\mathcal{X}}_{s}$ and every trajectory that starts from an unsafe state enters ${\mathcal{X}}_{s}$ in a finite number of steps and there remains?
In Fornasini & Valcher (2016) it has been provided a complete characterization of the problem solvability.
Proposition 12.
Given a BCN (4a) and the set ${\mathcal{X}}_{u}$ of unsafe states, the safe control problem is solvable if and only if
  • (i)
    for every ${\bf x}\in{\mathcal{X}}_{s}={\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}$ there exists ${\bf u}\in{\mathcal{L}}_{M}$ such that $L\ltimes{\bf u}\ltimes{\bf x}\in{\mathcal{X}}_{s}$;
  • (ii)
    the set ${\mathcal{X}}_{s}$ is reachable from every ${\bf x}\in{\mathcal{X}}_{u}$, which amounts to saying that for every ${\bf x}\in{\mathcal{X}}_{u}$ there exists $\bar{\bf x}\in{\mathcal{X}}_{s}$ such that $\bar{\bf x}$ is reachable from ${\bf x}$.
Moreover, if the safe control problem is solvable, then it is solvable by means of a state feedback law.
The data-driven version of the safe control problem is the following one.
Problem 2.
Given the data $({\bf x}_{d}$, ${\bf u}_{d})$ and the unsafe set ${\mathcal{X}}_{u}\subsetneq{\mathcal{L}}_{N}$, determine (if it exists) a state feedback control law ${\bf u}(t)=K{\bf x}(t)$, $K\in\mathcal{L}_{M\times N}$, that solves the safe control problem for every BCN compatible with the data.
When Problem 2 is solvable, the data $({\bf x}_{d}$, ${\bf u}_{d})$ will be called informative for safe control (with respect to the unsafe set ${\mathcal{X}}_{u}$). The problem solution relies on the tools developed to solve Problem 1. First of all, we need to identify from data if each state in ${\mathcal{X}}_{s}$ has (at least) one successor in ${\mathcal{X}}_{s}$. This amounts to verifying if for every $\delta^{j}_{N}\in{\mathcal{X}}_{s}$ there exists $k\in[1,T]$ and $\delta^{\ell}_{N}\in{\mathcal{X}}_{s}$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}.$ (15)
If this is the case we can set
${\bf u}_{j}^{*}\triangleq U_{p}\delta_{T}^{k}.$ (16)
If the previous check gives a positive answer, then condition (i) in Proposition 12 holds for every BCN in ${\mathcal{B}}_{d}$. At this point we need to check that condition (ii) holds for every BCN in ${\mathcal{B}}_{d}$. This is possible by extending Algorithm 2.
Algorithm 3: Reachability of ${\mathcal{X}}_{s}$.
Input: - The data matrices $X_{p},X_{f}$ and $U_{p}$;
           - The set ${\mathcal{X}}_{s}=\{\delta^{j_{1}}_{N},\delta^{j_{2}}_{N},\dots,\delta^{j_{d}}_{N}\}$.
Output: - Is ${\mathcal{X}}_{s}$ globally reachable for every BCN in ${\mathcal{B}}_{d}$?
               Yes/No.
             - The basin of attraction ${\mathcal{S}}^{*}\triangleq\bigcup_{{\bf x}\in{\mathcal{X}}_{s}}\mathcal{S}({\bf x})$ of
               ${\mathcal{X}}_{s}$.
Initialization: $({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
for $r\in[1,N]$, do
     if $\delta^{r}_{N}\in\mathcal{X}_{s}$, then
         ${\bf u}_{r}\leftarrow{\bf u}_{r}^{*}$ (see (16))
     end if
end for
Set $d=0$, ${\mathcal{S}}_{0}=\mathcal{X}_{s}$.
Iterative procedure:
if ${\mathcal{S}}_{d}\neq\emptyset$, then
  • $d\leftarrow d+1$
  • ${\mathcal{S}}_{d}=\emptyset$
  • for $k\in[1,T]$, do
            if
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix},\qquad\exists\ \delta^{q}_{N}\in{\mathcal{S}}_{d-1},$
    and $\delta^{r}_{N}\not\in\cup_{i=1}^{d}{\mathcal{S}}_{i}$, then
                   ${\mathcal{S}}_{d}\leftarrow{\mathcal{S}}_{d}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}\triangleq U_{p}\delta_{T}^{k}$
            end if
    end for
  • Go back to the Iterative procedure.
else set $\mathcal{S}^{*}=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$ and go to the Conclusion
end if
Conclusion: if $\mathcal{S}^{*}=\mathcal{L}_{N}$, then return (‘Yes’, $\mathcal{S}^{*}$)
                else, return (‘No’, $\mathcal{S}^{*}$)
                end if
Theorem 13.
Problem 2 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d})$ are informative for safe control with respect to the unsafe set ${\mathcal{X}}_{u}$) if and only if the following two conditions hold:
  • (i)
    for every $\delta_{N}^{j}\in{\mathcal{X}}_{s}$ there exist $k\in[1,T]$ and $\delta_{N}^{\ell}\in{\mathcal{X}}_{s}$ such that (15) holds;
  • (ii)
    ${\mathcal{S}^{*}}=\bigcup_{{\bf x}\in\mathcal{X}_{s}}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
Moreover, the generic $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by  ${\rm col}_{r}(K)={\bf u}_{r},\forall r\in[1,N],$ where ${\bf u}_{r}$ is determined through Algorithm 3.
Proof.
(Necessity). Assume, by contradiction, that condition (i) does not hold. This implies that there is at least one state in $\mathcal{X}_{s}$ whose immediate successors, as they can be deduced from data, are all in $\mathcal{X}_{u}$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which condition (i) in Proposition 12 does not hold. For that BCN the safe control problem is not solvable.
Assume now that (i) holds, but (ii) does not, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We want to construct a BCN in ${\mathcal{B}}_{d}$ for which one of the states in $\mathcal{X}_{u}$ cannot reach a safe state in a finite number of steps. This requires to impose all the state transitions that are not uniquely identified by the data. If there exists $\delta^{r}_{N}\in{\mathcal{X}}_{u}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, $\mathcal{X}_{s}$ is not reachable from $\delta^{r}_{N}\in{\mathcal{X}}_{u}$ and hence by Proposition 12, the safe control problem for such BCN in ${\mathcal{B}}_{d}$ is not solvable.
(Sufficiency). It is a direct consequence of Proposition 12 and subsequent comments.
The second part follows trivially from Algorithm 3. ∎
Example 14, below, explores on a small size network the results of this section and of the previous one pertaining data-driven state-feedback stabilization and safe control.
Example 14.
Consider a BCN (4a) with $N=7$ and $M=3$, described by the following matrix
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\big{[}\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ \delta_{7}^{5}\ \delta_{7}^{2}\ \delta_{7}^{7}\ \delta_{7}^{5}\ |$
$|\ \delta_{7}^{6}\ \delta_{7}^{1}\ \delta_{7}^{3}\ \delta_{7}^{2}\ \delta_{7}^{4}\ \delta_{7}^{5}\ \delta_{7}^{7}\ |\ \delta_{7}^{7}\ \delta_{7}^{6}\ \delta_{7}^{2}\ \delta_{7}^{3}\ \delta_{7}^{1}\ \delta_{7}^{6}\ \delta_{7}^{6}\big{]}.$
The equilibrium points of this BCN are $\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{6},\delta_{7}^{7}\}$. We perform an offline experiment in the time interval $[0,T]$, with $T=12$ and we collect the following data:
$X_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{7}^{1}&\delta_{7}^{7}&\delta_{7}^{7}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{4}&\delta_{7}^{3}&\delta_{7}^{3}&\delta_{7}^{2}\end{array}\right],$
$U_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}\end{array}\right],$
$X_{f}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{7}^{7}&\delta_{7}^{7}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{4}&\delta_{7}^{3}&\delta_{7}^{3}&\delta_{7}^{2}&\delta_{7}^{2}\end{array}\right].$
The BCNs compatible with these data can be generically represented by the following matrix $\tilde{L}$
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\big{[}\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ *\ *\ *\ *\ |$
$|\ \delta_{7}^{6}\ *\ \delta_{7}^{3}\ *\ *\ \delta_{7}^{5}\ \delta_{7}^{7}\ |\ \delta_{7}^{7}\ *\ *\ \delta_{7}^{3}\ \delta_{7}^{1}\ *\ \delta_{7}^{6}\big{]}.$
where the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{7}$). By applying Algorithm 1, we obtain that the set of the equilibrium points compatible with all the BCNs in $\mathcal{B}_{d}$ is $\mathcal{X}_{e}^{d}=\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{7}\}.$ We now check if there exists at least one state in $\mathcal{X}_{e}^{d}$ that is globally reachable. To this end, we follow the procedure outlined in Algorithm 2.
$\bullet$  For ${\bf x}_{e}^{1}\triangleq\delta_{7}^{2}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{2}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{3}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{4}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{5}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{6}=\{\delta_{7}^{7}\}$, and hence $\mathcal{S}({\bf x}_{e}^{1})={\mathcal{L}}_{7}$.
$\bullet$   For ${\bf x}_{e}^{2}\triangleq\delta_{7}^{3}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{3}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{4}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{4}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{5}=\{\delta_{7}^{7}\}$, $\mathcal{S}_{6}=\emptyset$, and hence $\mathcal{S}({\bf x}_{e}^{2})\neq{\mathcal{L}}_{7}$.
$\bullet$   For ${\bf x}_{e}^{2}\triangleq\delta_{7}^{7}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{7}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{4}=\emptyset$, and hence $\mathcal{S}({\bf x}_{e}^{2})\neq{\mathcal{L}}_{5}$.
Therefore, by Theorem 10, the data are informative for stabilization to ${\bf x}_{e}^{1}$ by state feedback. The state feedback matrix $K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}\in\mathcal{L}_{3\times 7}$ is obtained through Algorithm 2.
Assume now that we want to implement a state feedback safe control with respect to the unsafe set $\mathcal{X}_{u}=\{\delta_{7}^{3},\delta_{7}^{4},\delta_{7}^{7}\}$. We need to understand if the data are informative for safe control with respect to $\mathcal{X}_{u}$, or, equivalently, if conditions (i) and (ii) of Theorem 13 are satisfied. In this case $\mathcal{X}_{s}=\{\delta_{7}^{1},\delta_{7}^{2},\delta_{7}^{5},\delta_{7}^{6}\}$, and hence it is easy to check on data that for every state in $\mathcal{X}_{s}$ there is a transition ending in a state that is still inside $\mathcal{X}_{s}$, namely condition (i) of Theorem 13 holds. Now we apply Algorithm 3 to verify whether $\mathcal{X}_{s}$ is globally reachable in every BCN in ${\mathcal{B}}_{d}$. We obtain: $\mathcal{S}_{0}=\mathcal{X}_{s}$, $\mathcal{S}_{1}=\{\delta_{7}^{3},\delta_{7}^{7}\}$, and $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, and hence $\mathcal{S}^{*}=\mathcal{L}_{7}$. This means that also condition (ii) of Theorem 13 is satisfied, and thus the data are informative for safe control with respect to $\mathcal{X}_{u}$. Algorithm 3 provides the state feedback matrix $K=\begin{bmatrix}\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}$.
5 Data-driven output regulation
As before, we assume to collect some input and state data on a finite time interval $[0,T],T\in\mathbb{Z}_{+}$. However, this time we collect also output trajectories. Let ${\bf y}_{d}\triangleq\{{\bf y}_{d}(t)\}_{t=0}^{T-1}$ be the corresponding output sequence. We group the output data into a matrix analogous to the ones in (9)$\div$(11):
$Y_{p}\triangleq\begin{bmatrix}{\bf y}_{d}(0)&{\bf y}_{d}(1)&\dots&{\bf y}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{P\times T}.$ (22)
We extend the concept of compatibility of a BCN with the collected data to the case in which the BCN is described by (4), i.e., we include the output equation, and the collected data are $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$. The set of Boolean control networks (4) compatible with $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ is
$\!\!\!\!\mathcal{B}_{d}^{o}\!$$\triangleq\{(\tilde{L},\tilde{H}),\tilde{L}\in\mathcal{L}_{N\times NM},\tilde{H}\in\mathcal{L}_{P\times N}:\forall\ t\in[0,T-1]$
$\!\!\!{\bf x}_{d}(t+1)=\tilde{L}\ltimes{\bf u}_{d}(t)\ltimes{\bf x}_{d}(t),\ {\bf y}_{d}(t)=\tilde{H}{\bf x}_{d}(t)\}.\!\!\!$ (23)
Obviously, the pair $(L,H)$ corresponding to the BCN (4) that generated the data belongs to $\mathcal{B}_{d}^{o}$. Accordingly, the notion of informativity for identifiability has to be extended.
Definition 15.
The data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are called informative for identifiability of the BCN (4) if ${\mathcal{B}}_{d}^{o}\equiv\{(L,H)\}$.
Even if the definition is different, the characterization of the new notion of informativity for identifiability is still the one in Lemma 6: the matrix in (13) must be of full row rank. Necessity is obvious. On the other hand, if the matrix in (13) is of full row rank, then we deduce from the collected data every possible state transition that can occur in the network. But this implies that all possible states appear at least once in $X_{p}$ (and consequently also the corresponding output in $Y_{p}$). This allows to uniquely identify also the matrix $H$ in (4b), and hence the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the whole BCN (4).
In the sequel, we assume again that the data are not informative for the identifiability of the BCN.
Problem 3.
Given the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$, and some desired value ${\bf y}^{*}\in{\mathcal{L}}_{P}$ for the output in (4b), determine (if possible) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that, by applying ${\bf u}(t)=K{\bf x}(t)$, for every initial condition ${\bf x}(0)$, the output trajectories of all the BCNs compatible with the data satisfy ${\bf y}(t)={\bf y}^{*}$ for every $t\geq\tau,\exists\tau\in\mathbb{Z}_{+}$.
When Problem 3 is solvable, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback. In order to solve this problem, we introduce some definitions and results, taken from the model-based literature on BNs and BCNs (see, e.g., Fornasini & Valcher (2013a, 2014)).
Definition 16.
Given a BCN (4), the output regulation problem to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback is said to be solvable if there exist $K\in\mathcal{L}_{M\times N}$ and $\tau\in\mathbb{Z}_{+}$ such that for every ${\bf x}(0)\in\mathcal{L}_{N}$, ${\bf y}(t)={\bf y}^{*},\forall t\geq\tau$, when ${\bf u}(t)=K{\bf x}(t).$
The first step of the model-based solution to this problem consists in finding all the states that generate the desired output ${\bf y}^{*}$, namely the states belonging to the set $\mathcal{X}({\bf y}^{*})\triangleq\{\delta_{N}^{j}:H\delta_{N}^{j}={\bf y}^{*}\}$. Once we have determined this set, we have to ensure that, under some state feedback law, the state evolution of the resulting BN (8) is constrained within $\mathcal{X}({\bf y}^{*})$, from some time $\tau\in\mathbb{Z}_{+}$ onwards, which imposes that $\mathcal{X}({\bf y}^{*})$ contains either an equilibrium point of the BN or a limit cycle (whose definition is given below).
Definition 17.
Fornasini & Valcher (2013b) Given a BN as in (3), an ordered sequence of distinct logical vectors $(\delta_{N}^{i_{1}},\delta_{N}^{i_{2}},\dots,\delta_{N}^{i_{k}})$ is a limit cycle $\mathcal{C}$ if, taken ${\bf x}(0)=\delta_{N}^{i_{l}}$ for some $l\in[1,k]$, the corresponding state evolution ${\bf x}(t)$ is periodic of period $k$ and, for every $t\in\mathbb{Z}_{+}$, ${\bf x}(t)=\delta_{N}^{i_{j}}$, with $j\in[1,k]$ satisfying $j=(t+l){\rm mod}k$.
Note that an equilibrium point is a limit cycle of unitary length. We denote by $\mathcal{C}({\bf y}^{*})$ the set of all limit cycles contained in $\mathcal{X}({\bf y}^{*})$. Once we have ascertained that $\mathcal{C}({\bf y}^{*})\neq\emptyset$, the final step consists in verifying that it is globally reachable, meaning that the union of the basins of attraction of the states in $\mathcal{C}({\bf y}^{*})$ covers the whole state space $\mathcal{L}_{N}$.
In the following proposition we state necessary and sufficient conditions for solving the output regulation problem by state feedback in a model-based set-up.
Proposition 18.
Fornasini & Valcher (2014) Given a BCN as in (4) and a desired output value ${\bf y}^{*}\in{\mathcal{L}}_{P}$, the problem of output regulation to ${\bf y}^{*}$ by state feedback is solvable if and only if
  • (i)
    $\mathcal{C}({\bf y}^{*})\neq\emptyset$;
  • (ii)
    $\mathcal{C}({\bf y}^{*})$ is globally reachable, i.e., $\cup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}{\mathcal{S}}({\bf x})\!=\!{\mathcal{L}}_{N}$.
We can now provide the solution to Problem 3. The first step consists in finding, from data, the set of states that generate the desired output ${\bf y}^{*}$ in all BCNs in ${\mathcal{B}}_{d}^{o}$, i.e.:
$\mathcal{X}^{d}({\bf y}^{*})\!=\!\left\{\delta_{N}^{j}\in{\mathcal{L}}_{N}:\exists k\in[1,T]\ \text{s.t.}\begin{bmatrix}X_{p}\cr Y_{p}\end{bmatrix}\delta_{N}^{k}=\!\begin{bmatrix}\delta_{N}^{j}\cr{\bf y}^{*}\end{bmatrix}\right\}\!.$
To identify this set, we can proceed as follows.
Algorithm 4: Construction of $\mathcal{X}^{d}({\bf y}^{*})$
Input: - The data matrices $X_{p}$ and $Y_{p}$;
           - ${\bf y}^{*}\in\mathcal{L}_{P}$.
Output: The set $\mathcal{X}^{d}({\bf y}^{*})$ of all states of the BCNs in ${\mathcal{B}}_{d}^{o}$ that generate the desired output ${\bf y}^{*}$.
Initialization:
Set $\mathcal{X}^{d}({\bf y}^{*})=\emptyset$.
Procedure:
  • Compute
    $\left[\begin{array}[]{c|c}I_{N}&\mathbbold{0}_{N\times P}\cr\hline\cr\mathbbold{0}_{1\times N}&({\bf y}^{*})^{\top}\end{array}\right]\left[\begin{array}[]{c}X_{p}\cr\hline\cr Y_{p}\end{array}\right]=\left[\begin{array}[]{c}X_{p}\cr\hline\cr{\bf v}^{\top}\end{array}\right]$
    Note that ${\bf v}^{\top}=({\bf y}^{*})^{\top}Y_{p}\in\mathcal{B}^{T}$, and $[{\bf v}]_{k}\neq 0$ if and only if ${\rm col}_{k}(X_{p})\in\mathcal{X}^{d}({\bf y}^{*})$.
  • Define $\mathcal{K}\triangleq\{k_{1},k_{2},\dots,k_{y}\}=$nonzero(${\bf v}$)
  • if $\mathcal{K}\neq\emptyset$, then
         for $k\in\mathcal{K}$, do
            if $\delta_{N}^{k}\notin\mathcal{X}^{d}({\bf y}^{*})$, then
                $\mathcal{X}^{d}({\bf y}^{*})\leftarrow\mathcal{X}^{d}({\bf y}^{*})\cup\{\delta_{N}^{k}\}$
            end if
         end for
    end if
Conclusion: Return $\mathcal{X}^{d}({\bf y}^{*})$. ∎
At this point, we need to verify if $\mathcal{X}^{d}({\bf y}^{*})$ contains some limit cycles. It is worth highlighting that to determine limit cycles of unitary length we can use Algorithm 1. However, the next algorithm is able to also identify equilibrium points, being them a special case of limit cycles. To this end, let us define the data matrix
$X\!\triangleq\!\begin{bmatrix}{\bf x}_{d}(0)&{\bf x}_{d}(1)&\dots&{\bf x}_{d}(T-1)&{\bf x}_{d}(T)\end{bmatrix}\in\mathcal{L}_{N\times(T+1)},$
containing all the logical vectors of the sequence ${\bf x}_{d}$. The first step consists in determining all the transitions appearing in the matrix $X$ involving (both initial and final) states that belong to $\mathcal{X}^{d}({\bf y}^{*})$. Based on these transitions, we then construct a digraph, whose set of nodes is $\mathcal{X}^{d}({\bf y}^{*})$, and with edges corresponding to the transitions identified from the data. We call $\tilde{L}_{tot}$ the Boolean (but not necessarily logical) matrix associated to such a digraph, since we are accounting for transitions associated to different inputs. The matrix $\tilde{L}_{tot}$ represents the data-based estimate of the principal submatrix of $L_{tot}$ in (7) obtained by selecting the rows and the columns corresponding to the states in $\mathcal{X}^{d}({\bf y}^{*})$. Indeed, as the data are not informative for identifiability of the BCN, the unitary entries of $\tilde{L}_{tot}$ in general are only a subset of those of $L_{tot}$ for the original BCN (and for every BCN in ${\mathcal{B}}_{d}^{0}$). We now exploit the well-know Johnson’s Algorithm Johnson (1975) to identify all limit cycles in ${\mathcal{D}}(\tilde{L}_{tot})$.
Algorithm 5333The proposed algorithm is tailored for a sequence of data collected in a single experiment. If the performed experiments are more than one, it must be applied separately to each experiment, and then the results must be integrated. : Limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$
Input: - The data matrix $X$;
           - The set $\mathcal{X}^{d}({\bf y}^{*})$.
Output: The set $\mathcal{C}^{d}({\bf y}^{*})$ of all limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$.
Initialization:
Set:
  • $P\triangleq\begin{bmatrix}\delta_{N}^{i_{1}}&\delta_{N}^{i_{2}}&\dots&\delta_{N}^{i_{s}}\end{bmatrix}\in{\mathcal{L}}_{N\times s},$ where $s$ is the cardinality of $\mathcal{X}^{d}({\bf y}^{*})$ and $\mathcal{I}\triangleq\{i_{1},i_{2},\dots,i_{s}\}\subseteq[1,N]$ is the set of indices of the states in $\mathcal{X}^{d}({\bf y}^{*})$ (taken in an arbitrary order);
  • $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$;
  • $\mathcal{D}(\tilde{L}_{tot})\triangleq(\tilde{\mathcal{V}},\tilde{\mathcal{E}})$, with $\tilde{\mathcal{V}}=\mathcal{X}^{d}({\bf y}^{*})$ and $\tilde{\mathcal{E}}=\emptyset$.
Procedure:
if $\mathcal{I}\neq\emptyset$, then
  • Compute $X^{*}\triangleq P^{\top}X\in\mathcal{L}_{s\times(T+1)}$
  • for $(i,j)\in[1,s]\times[1,T+1]$, do
          if $\exists k\in[1,T]$ s.t. $[X^{*}]_{ik}[X^{*}]_{j,k+1}\neq 0$, then
              ~~$\tilde{\mathcal{E}}\leftarrow\tilde{\mathcal{E}}\cup\{(i,j)\}$
          end if
    end for
  • Apply Johnson’s Algorithm in Johnson (1975) to find all cycles in $\mathcal{D}(\tilde{L}_{tot})$, say $\tilde{\mathcal{C}}_{1},\dots,\tilde{\mathcal{C}}_{n_{c}}$, $n_{c}\in\mathbb{Z}_{+}$
  • Remap all the nodes of $\mathcal{D}(\tilde{L}_{tot})$ to the corresponding nodes in $[1,N]$, by using the bijective correspondence given by the matrix $P$, i.e., $1\leftrightarrow i_{1},\dots,s\leftrightarrow i_{s}$
  • Define $\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}$ the remapped limit cycles
  • $\mathcal{C}^{d}({\bf y}^{*})=\{\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}\}$
else Go to the Conclusion
end if
Conclusion: Return $\mathcal{C}^{d}({\bf y}^{*})$. ∎
For each edge $(i,j)$ belonging to some limit cycle ${\mathcal{C}}_{l}$ in $\mathcal{C}^{d}({\bf y}^{*})$ it is also possible to determine one of the inputs related to such transition within the limit cycle, say ${\bf u}_{i}^{l}$. Indeed, if there exists an edge $(i,j)$, it means that the transition $i\to j$ must be captured by the collected data, or, equivalently, there exists $k\in[1,T]$ such that $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{i}\cr\delta_{N}^{j}\end{bmatrix}.$ Therefore, the input associated to this transition within ${\mathcal{C}}_{l}$ is ${\bf u}_{i}^{l}\triangleq U_{p}\delta_{T}^{k}.$ At this point it only remains to establish if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. This is possible by applying (a slightly amended version of) Algorithm 3.
We can now provide the solution to Problem 3.
Theorem 19.
Problem 3 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback) if and only if the following two conditions hold:
  • (i)
    $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$;
  • (ii)
    ${\mathcal{S}^{*}_{o}}\triangleq\bigcup_{{\bf x}\in\mathcal{C}^{d}({\bf y}^{*})}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
Moreover, the generic $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by ${\rm col}_{r}(K)={\bf u}_{r},\forall r\in[1,N],$ where ${\bf u}_{r}$ is determined by Algorithm 3.
Proof.
(Necessity). Assume, by contradiction, that condition (i) does not hold, i.e., $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\emptyset$, and hence, by Proposition 18, for that BCN the output regulation to ${\bf y}^{*}$ by state feedback is not possible.
Assume now that $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$, but (ii) does not hold, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We can construct a BCN in ${\mathcal{B}}_{d}^{o}$ for which $\mathcal{C}({\bf y}^{*})=\mathcal{C}^{d}({\bf y}^{*})$ and $\bigcup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}\mathcal{S}({\bf x})={\mathcal{S}^{*}_{o}}\neq\mathcal{L}_{N}$. This requires to impose all the state transitions, equivalently the columns of $\tilde{L}$, that are not uniquely identified by the data (and possibly also the corresponding output). If there exists $\delta^{r}_{N}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ and $H\delta^{r}_{N}\neq{\bf y}^{*}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, by Proposition 18, for each BCN obtained in this way the output regulation to ${\bf y}^{*}$ by state feedback is not possible.
(Sufficiency). It is a direct consequence of Proposition 18 and following comments.
The final part follows trivially from Algorithm 3. ∎
Example 20.
Consider a BCN (4) with $N=6$, $M=3$, and $P=2$, described by the following matrices
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\left[\begin{array}[]{c|}\delta_{6}^{2}\ \delta_{6}^{4}\ \delta_{6}^{3}\ \delta_{6}^{3}\ \delta_{6}^{6}\ \delta_{6}^{5}\end{array}\right.$
$\left.\begin{array}[]{|c|c}\delta_{6}^{1}\ \delta_{6}^{5}\ \delta_{6}^{2}\ \delta_{6}^{2}\ \delta_{6}^{6}\ \delta_{6}^{1}&\delta_{6}^{5}\ \delta_{6}^{1}\ \delta_{6}^{4}\ \delta_{6}^{5}\ \delta_{6}^{4}\ \delta_{6}^{6}\end{array}\right],$
$H$ $=$ $\left[\begin{array}[]{c}\delta_{2}^{1}\ \delta_{2}^{2}\ \delta_{2}^{2}\ \delta_{2}^{2}\ \delta_{2}^{1}\ \delta_{2}^{1}\end{array}\right].$
Assume that we want to regulate the output of the network to the value ${\bf y}^{*}=\delta_{2}^{2}$. We perform an offline experiment in the time interval $[0,9]$, and collect the data:
$X$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c}\delta_{6}^{6}&\delta_{6}^{6}&\delta_{6}^{1}&\delta_{6}^{2}&\delta_{6}^{5}&\delta_{6}^{4}&\delta_{6}^{2}&\delta_{6}^{4}&\delta_{6}^{3}&\delta_{6}^{3}\end{array}\right],$
$U_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c}\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}\end{array}\right],$
$Y_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c}\delta_{2}^{1}&\delta_{2}^{1}&\delta_{2}^{1}&\delta_{2}^{2}&\delta_{2}^{1}&\delta_{2}^{2}&\delta_{2}^{2}&\delta_{2}^{2}&\delta_{2}^{2}\end{array}\right].$
We deduce $\mathcal{X}^{d}({\bf y}^{*})=\{\delta_{6}^{2},\delta_{6}^{3},\delta_{6}^{4}\}.$ The BCNs compatible with these data can be represented by the following matrix
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\left[\begin{array}[]{c|}\delta_{6}^{2}\ \delta_{6}^{4}\ \delta_{6}^{3}\ \delta_{6}^{3}\ *\ *\end{array}\right.$
$\left.\begin{array}[]{|c|c}*\ \delta_{6}^{5}\ *\ \delta_{6}^{2}\ *\ \delta_{6}^{1}&*\ *\ *\ *\ \delta_{6}^{4}\ \delta_{6}^{6}\end{array}\right],$
where, again, the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{6}$). As $X_{p}$ is of full row rank, the matrix $H$ can be uniquely recovered. Hence, by applying Algorithm 4, we obtain that the set of all the limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$ is $\mathcal{C}^{d}({\bf y}^{*})=\{\delta_{6}^{3},\delta_{6}^{2}\leftrightarrow\delta_{6}^{4}\},$ where $\delta_{6}^{2}\leftrightarrow\delta_{6}^{4}$ is the limit cycle of length $2$ involving $\delta_{6}^{2}$ and $\delta_{6}^{4}$. We now check if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. By following the procedure outlined in Algorithm 3, we obtain $\mathcal{S}_{o}^{*}={\mathcal{L}}_{6}$. Thus, by Theorem 19, the data are informative for output regulation to ${\bf y}^{*}$ by state feedback. The feedback matrix $K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}\end{bmatrix}$ is obtained via Algorithm 3.
6 Conclusion
In this paper we have investigated how the recently introduced data informativity approach van Waarde et al. (2020) can be adapted and extended to solve fundamental control problems in Boolean control networks. More specifically, we focused on the feedback stabilization, safe control and output regulation problems, and we provided necessary and sufficient conditions, based only on previously collected data (supposed not to be informative for identification), for their solvability. Moreover, we also obtained from data the feedback matrices expressions. We validated our results by means of examples.