### v2

# Data-based control of Logical Networks

Abstract
In recent years, data-driven approaches have become increasingly pervasive across all areas of control engineering. However, the applications of data-based techniques to Boolean control networks (BCNs) are still very limited. In this paper we aim to fill this gap, by exploring the possibility of evaluating some basic features, i.e., reachability and equilibria, and of solving two fundamental control problems, i.e., safe control and output regulation, for a BCN, leveraging only a limited amount of data generated by the network, without knowing or identifying its model.
keywords:
Boolean control networks; safe control; output regulation.
thanks: Corresponding Author.
1 Introduction
Boolean networks (BNs) are discrete-time (autonomous) dynamical systems whose describing variables are constrained to take values in the Boolean set. When their evolution is controlled by an external signal (an input), whose entries in turn take Boolean values, they are called Boolean control networks (BCNs). The interest in this class of logical systems originated in the 1960s, when Stuart Kauffman introduced them to model and analyze gene regulatory networks Kauffman (1969), and even if this may still be regarded as their most successful application area, they have been adopted in a wide range of fields, including systems biology, computer science, game theory, smart cities and network theory (see, e.g., Bornholdt (2008); Cheng (2014); Green et al. (2007); Kabir et al. (2014); Schreiber & Valcher (2022)).
The algebraic approach to BNs and BCNs, proposed by Cheng et al. Cheng (2009); Cheng & Qi (2010a, b); Cheng et al. (2011), and deeply relying on the concept of semi-tensor product (STP) Cheng (2001), has made it possible to represent BNs and BCNs by means of discrete-time state space models in which the state, input and output are logical vectors, and the describing matrices are logical matrices. In this way, concepts and tools developed for linear time-invariant state space models have been adjusted to formalize and solve control problems for BNs and BCNs (see, e.g., Fornasini & Valcher (2014); Li et al. (2013); Zhang & Zhang (2016)).
Despite the simplifying “on/off" (“active/inactive", “high/low") logic at the basis of BCNs, the size and complexity of the physical systems they model (e.g., gene regulatory networks) make it difficult to derive an accurate model of the network, as well as to obtain it by means of identification techniques. On the other hand, in recent times direct data-driven methods have gained interest in the control community, given their attractive feature of bypassing the identification phase, to directly design the problem solution based on raw data. Yet, the use of partial data to solve control problems for logical networks is still at an early stage. The first attempt in this direction has been made in Leifeld et al. (2018), where a one-step data-driven approach to solve the output stabilization problem is proposed. More specifically, the authors in Leifeld et al. (2018) first reconstruct from the available data the output prediction matrix that maps the current input/output pair to the set of all possible next outputs, and then, based on this map, design an output feedback law to stabilize the output of the BCN to a desired value. However, in general, the set of all possible next outputs contains more than one element and hence the output prediction is possibly non-deterministic. To deal with this uncertainty, the authors draw inspiration from results on the stabilization of Probabilistic Boolean networks (PBNs).
The authors in Li et al. (2025), instead, exploit the informativity framework, introduced by van Waarde and coauthors in van Waarde et al. (2020) (see, also, van Waarde et al. (2023)), to investigate BCN control problems in a data-driven set-up. The informativity framework is based on the idea that, when data are not informative enough to uniquely identify the system, the best we can do is impose that all the systems compatible with the collected data (among which there is surely also our target system) possess the desired property or behave in a certain way. In Li et al. (2025) fundamental system properties such as controllability and stabilizability are investigated, and the problems of state feedback stabilization and guaranteed cost control are addressed for BCNs. For both problems, necessary and sufficient conditions under which the data are informative for control design are provided.
In this paper, we also exploit this approach to address the limitations arising from having only a limited amount of data available. More specifically, we solve two fundamental control problems, namely safe control and output regulation of a BCN via state feedback, by solving them for all logical systems compatible with the available data. This automatically guarantees that the problem is solved for the actual BCN, even under partial information. We provide necessary and sufficient conditions on the collected data to drive the evolution of a target BCN towards a safe set, and to set its output to a desired value, by means of a state feedback. Moreover, for each of these problems, we propose a procedure to design such a feedback matrix based on the collected data. In order to solve these problems, we preliminarily investigate under what conditions the reachability of a set of states and the existence of equilibrium points can be inferred, for all the BCNs compatible with the data, directly from the data themselves, since these notions prove to be instrumental in constructing the solutions.
The paper is organized as follows. In Section 2 we briefly introduce Boolean networks and Boolean control networks. In Section 3 we explain the data collection process and we provide the notion of informativity for identifiability, along with its characterization. Section 4 contains an informativity-based characterization of the reachability property for a BCN. Section 5 reviews the notion of equilibrium points in BCNs and provides a systematic approach to identifying all equilibrium points consistent with the data. In Section 6 we formalize and solve the safe control problem for a BCN using only finite sequences of data. In Section 7 the data-driven output regulation problem by state feedback for a BCN is investigated. Section 8 concludes the paper.
Notation. Given two nonnegative integers $k$ and $n$, with $k\leq n$, we denote by $[k,n]$ the set of integers $\{k,k+1,\dots,n\}$. We consider Boolean vectors and matrices whose entries take values in $\mathcal{B}\triangleq\{0,1\}$, equipped with the usual logical operations, namely sum (OR) $\vee$, product (AND) $\land$, and negation (NOT) $\neg$. We denote by $\delta_{k}^{i}$ the $k$-dimensional $i$-th canonical vector, by $\mathcal{L}_{k}$ the set of all $k$-dimensional canonical vectors, and by $\mathcal{L}_{k\times q}\subset\mathcal{B}^{k\times q}$ the set of all $k\times q$ logical matrices, namely matrices whose $q$ columns are canonical vectors of size $k$. Any matrix $L\in\mathcal{L}_{k\times q}$ can be described as $L=[\delta_{k}^{i_{1}}\ \delta_{k}^{i_{2}}\ \dots\ \delta_{k}^{i_{q}}]$, for some indices $i_{1},i_{2},\dots,i_{q}\in[1,k]$. $I_{n}$ denotes the $n$-dimensional identity matrix. The $k$-dimensional vector with all unitary (all zero) entries is denoted by $\mathbbold{1}_{k}$ (by $\mathbbold{0}_{k}$), and the $k\times l$ matrix with all zero entries by $\mathbbold{0}_{k\times l}$. When clear from the context or unnecessary, the suffixes are omitted. The $(i,j)$-th entry of a matrix $M$ is denoted by $M]_{ij}$, and the $j$-th entry of a vector $\bf v$} by ${\bf v}]_{j}$. The $j$-th column of matrix $M$ is denoted by ${\rm col}_{j}(M)$. We denote by $\odot$ the Hadamard (or entry-wise) product. Given two Boolean matrices $M_{1}$ and $M_{2}$ of the same size, condition $M_{1}\leq M_{2}$ means that $M_{1}]_{ij}\leq[M_{2}]_{ij}$ for every $i,j$. Given a vector ${\bf v}\in{\mathbb{R}}^{k}$, we denote by ${\rm diag}({\bf v}^{\top})$ the $k\times k$ diagonal matrix whose $(i,i)$-th entry is ${\bf v}]_{i}$. A Boolean (in particular, a logical) square matrix $L$ of size $k$ is reducible if there exists a permutation matrix $P\in\mathcal{L}_{k\times k}$ such that
$P^{\top}LP=\begin{bmatrix}L_{11}&L_{12}\cr\mathbbold{0}&L_{22}\end{bmatrix}$,
where $L_{11}$ and $L_{22}$ are square Boolean (logical) matrices. Otherwise, $L$ is called irreducible. $L$ is irreducible if and only if $I_{k}\vee L\vee\dots\vee L^{k-1}$ has all unitary entries.
There is a bijective correspondence between Boolean variables $X\in\mathcal{B}$ and vectors ${\bf x}\in\mathcal{L}_{2}$, defined by ${\bf x}=[X\ \neg X]^{\top}$. Given $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{p\times q}$, their (left) semi-tensor product ($\ltimes$) (see Cheng (2001)) is
$A\ltimes B\triangleq(A\otimes I_{l/n})(B\otimes I_{l/p}),\quad l\triangleq{\rm l.c.m.}\{n,p\},$
where the symbol $\otimes$ denotes the Kronecker product. For the properties of the semi-tensor product, we refer to Cheng et al. (2011).
Given two matrices $C\in{\mathcal{B}}^{m\times n}$ and $D\in{\mathcal{B}}^{p\times n}$ with the same number of columns, we denote by * their Khatri-Rao product, which is defined as a column-wise semi-tensor product, or, equivalently, a column-wise Kronecker product, i.e.,
$C*D$ $=$ $\left[\begin{array}[]{c|c|c}{\rm col}_{1}(C)\ltimes{\rm col}_{1}(D)&\dots&{\rm col}_{n}(C)\ltimes{\rm col}_{n}(D)\end{array}\right]$
$=$ $\left[\begin{array}[]{c|c|c}{\rm col}_{1}(C)\otimes{\rm col}_{1}(D)&\dots&{\rm col}_{n}(C)\otimes{\rm col}_{n}(D)\end{array}\right].$
By exploiting the semi-tensor product, it is possible to extend the previously mentioned bijective correspondence to a bijective correspondence between $\mathcal{B}^{n}$ and $\mathcal{L}_{2^{n}}$, as follows. Given a vector $X=[X_{1}\ X_{2}\ \dots\ X_{n}]^{\top}\in\mathcal{B}^{n}$, set
${\bf x}\triangleq\begin{bmatrix}X_{1}\\ \neg X_{1}\end{bmatrix}\ltimes\begin{bmatrix}X_{2}\\ \neg X_{2}\end{bmatrix}\ltimes\dots\ltimes\begin{bmatrix}X_{n}\\ \neg X_{n}\end{bmatrix}$.
It is possible to associate with a Boolean (in particular, a logical) matrix $L\in{\mathcal{B}}^{k\times k}$ a directed graph (or, digraph), $\mathcal{D}(L)=(\mathcal{V},\mathcal{E})$, where ${\mathcal{V}}=[1,k]$ is the set of nodes and $\mathcal{E}$} is the set of edges. An edge $(j,l)$, from $j$ to $l$, belongs to $\mathcal{E}$ if and only if $L]_{lj}=1$. A sequence $j_{1}\to j_{2}\to\dots\to j_{r}\to j_{r+1}$ in $\mathcal{D}(L)$ is a path of length $r$ from $j_{1}$ to $j_{r+1}$ if $(j_{1},j_{2}),\dots,(j_{r},j_{r+1})$ are edges in $\mathcal{E}$. A closed path is a cycle. A node $j^{*}\in{\mathcal{V}}$ is said to be globally reachable if there exists a path to $j^{*}$ from any other node in the network. A digraph $\mathcal{D}(L)$ is strongly connected if all its nodes are globally reachable, and this is the case if and only if $L$ is irreducible.
2 Introduction to Boolean control networks
A Boolean control network (BCN) is a logical system described by the following equations
$X(t+1)$ $=$ $f(X(t),U(t)),$ (3a)
$Y(t)$ $=$ $h(X(t)),\qquad\qquad t\in\mathbb{Z}_{+},$ (3b)
where $X(t)\in\mathcal{B}^{n}$ is the $n$-dimensional state variable, $U(t)\in\mathcal{B}^{m}$ is the $m$-dimensional input, and $Y(t)\in\mathcal{B}^{p}$ is the $p$-dimensional output at time $t$. $f$ and $h$ are logical functions, i.e., $f:\mathcal{B}^{n}\times\mathcal{B}^{m}\to\mathcal{B}^{n}$, while $h:\mathcal{B}^{n}\to\mathcal{B}^{p}$. If the logical system is autonomous, namely no input acts on it, then the BCN becomes a Boolean network (BN) and equation (3a) becomes
$X(t+1)=f(X(t)),\quad t\in\mathbb{Z}_{+},$ (4)
where $f:\mathcal{B}^{n}\to\mathcal{B}^{n}$ is a logical function. By exploiting the bijective correspondence between Boolean and logical vectors, the BN (4) can be equivalently expressed via its algebraic representation Cheng et al. (2011) as
${\bf x}(t+1)=L{\bf x}(t),\quad t\in\mathbb{Z}_{+},$ (5)
where ${\bf x}(t)\in\mathcal{L}_{N}$ and $L\in{\mathcal{L}}_{N\times N}$, with $N\triangleq 2^{n}$. Similarly, the algebraic representation of the BCN (3) is
${\bf x}(t+1)$ $=$ $L\ltimes{\bf u}(t)\ltimes{\bf x}(t),\quad t\in\mathbb{Z}_{+},$ (6a)
${\bf y}(t)$ $=$ $H{\bf x}(t),$ (6b)
where ${\bf x}(t)\in\mathcal{L}_{N}$, ${\bf u}(t)\in\mathcal{L}_{M}$, ${\bf y}(t)\in\mathcal{L}_{P}$, $L\in\mathcal{L}_{N\times NM}$ and $H\in{\mathcal{L}}_{P\times N}$, with $N\triangleq 2^{n}$, $M\triangleq 2^{m}$ and $P\triangleq 2^{p}$. It is worth remarking that model (6) can be used to represent any state-space model whose state, input and output variables take values in sets of finite cardinalities $N$, $M$ and $P$, respectively, and hence all the subsequent analysis holds for generic nonnegative integer values of $N$, $M$ and $P$, not necessarily powers of $2$. The matrix $L\in\mathcal{L}_{N\times NM}$, whose columns are canonical vectors of size $N$, can be divided into $M$ square blocks of size $N$ as follows
$L=\begin{bmatrix}L_{1}&\vline&L_{2}&\vline&\dots&\vline&L_{M}\end{bmatrix}$, (7)
where each block $L_{i}\in\mathcal{L}_{N\times N},i\in[1,M]$, describes the behavior of the $i$-th subsystem of the BCN, i.e., the Boolean network obtained when ${\bf u}(t)=\delta_{M}^{i},\forall t\in\mathbb{Z}_{+}$:
${\bf x}(t+1)=L_{i}{\bf x}(t),\quad t\in\mathbb{Z}_{+}.$ (8)
3 Data collection process and informativity for identifiability of BCNs
We assume to have performed some offline experiments [^1] It is worth noticing that the number and the lengths of the experiments will play no role in the following analysis. However, a higher number of experiments may allow to obtain information that could not be collected in a single one., say $r\geq 1$, during which we have collected state/input/output data from the BCN (6) on finite time intervals $0,T_{i}],T_{i}\in\mathbb{Z}_{+},i\in[1,r]$. We define the vector sequences ${\bf x}_{d}^{i}\triangleq\{{\bf x}_{d}^{i}(t)\}_{t=0}^{T_{i}}$, ${\bf u}_{d}^{i}\triangleq\{{\bf u}_{d}^{i}(t)\}_{t=0}^{T_{i}-1},$ and ${\bf y}_{d}^{i}\triangleq\{{\bf y}_{d}^{i}(t)\}_{t=0}^{T_{i}-1},i\in[1,r]$, and accordingly ${\bf x}_{d}\triangleq\{{\bf x}_{d}^{i}\}_{i=1}^{r}$, ${\bf u}_{d}\triangleq\{{\bf u}_{d}^{i}\}_{i=1}^{r}$ and ${\bf y}_{d}\triangleq\{{\bf y}_{d}^{i}\}_{i=1}^{r}$. We rearrange the data collected during the $r$ experiments into the following logical matrices with $T\triangleq\sum_{i=1}^{r}{T_{i}}$ columns:
$X_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf x}_{d}^{1}(0)&\dots&{\bf x}_{d}^{1}(T_{1}-1)&{\bf x}_{d}^{2}(0)&\dots&{\bf x}_{d}^{r}(T_{r}-1)\end{array}\right],$
$X_{f}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf x}_{d}^{1}(1)&\dots&{\bf x}_{d}^{1}(T_{1})&{\bf x}_{d}^{2}(1)&\dots&{\bf x}_{d}^{r}(T_{r})\end{array}\right],$
$U_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf u}_{d}^{1}(0)&\dots&{\bf u}_{d}^{1}(T_{1}-1)&{\bf u}_{d}^{2}(0)&\dots&{\bf u}_{d}^{r}(T_{r}-1)\end{array}\right],$
$Y_{p}$ $\triangleq$ $\left[\begin{array}[]{ccc|ccc}{\bf y}_{d}^{1}(0)&\dots&{\bf y}_{d}^{1}(T_{1}-1)&{\bf y}_{d}^{2}(0)&\dots&{\bf y}_{d}^{r}(T_{r}-1)\end{array}\right],$
where the subscripts $p$ and $f$ stand for past and future, respectively.
As every BCN (6) bijectively corresponds to its describing matrices, we define the set of Boolean control networks (6) compatible with the data $({\bf x}_{d}$, ${\bf u}_{d},{\bf y}_{d})$ as
$\!\!\!\!\mathcal{B}_{d}\!$ $\triangleq\big\{(\tilde{L},\tilde{H})\in\mathcal{L}_{N\times NM}\times\mathcal{L}_{P\times N}:$
$\!\!\!{\bf x}_{d}^{i}(t+1)=\tilde{L}\ltimes{\bf u}_{d}^{i}(t)\ltimes{\bf x}_{d}^{i}(t),\ {\bf y}_{d}^{i}(t)=\tilde{H}{\bf x}_{d}^{i}(t),$
$\!\!\!\forall t\in[0,T_{i}-1],\forall i\in[1,r]\big\}.\!\!\!$ (13)
Obviously, the pair $(L,H)$ corresponding to the BCN (6) that generated the data belongs to $\mathcal{B}_{d}$. However, ${\mathcal{B}}_{d}$ may also include additional matrix pairs when the available data do not uniquely determine $(L,H)$. To this end, we introduce the notion of informativity for identifiability [^2] As a matter of fact, what we call informativity for identifiability is referred to as informativity for model reconstructability in Definition 3.3 of Li et al. (2025). , that extends Definition 3.3 in Li et al. (2025) to BCNs as in (6), which also include an output equation.
Definition 1.
The data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are called informative for identifiability of the BCN (6) if ${\mathcal{B}}_{d}\equiv\{(L,H)\}$.
Despite the presence of the output equation, the characterization of informativity for identifiability is the same as the one given in Proposition 3.4 of Li et al. (2025) for BCNs described only by (6a). Lemma 2 proves such equivalence and provides an additional equivalent condition.
Lemma 2.
The following facts are equivalent.
  • i)
    The data $({\bf x}_{d}$, ${\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the BCN (6).
  • ii)
    For every $(i,j)\in[1,M]\times[1,N]$, there exists $k\in[1,T]$ such that
    $\begin{bmatrix}X_{p}\cr U_{p}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{M}^{i}\end{bmatrix}$. (14)
  • iii)
    The logical matrix $U_{p}*X_{p}$ has no zero rows (equivalently, is of full row rank).
Proof.
i) $\Rightarrow$ iii) If i) holds, then the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of equation (6a), and hence, by Proposition 3.4 in Li et al. (2025), we can claim that iii) holds.
iii) $\Rightarrow$ i) If condition iii) holds, by Proposition 3.4 in Li et al. (2025), we can uniquely identify the matrix $L$. On the other hand, if the matrix $U_{p}*X_{p}$ is of full row rank, then $X_{p}$ is of full row rank, in turn. This means that all possible states appear at least once in $X_{p}$, and consequently also the corresponding output in $Y_{p}$. This allows to uniquely identify also the matrix $H$ in (6b), and hence the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the whole BCN (6).
ii) $\Leftrightarrow$ iii) is straightforward. ∎
Remark 3.
If the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability and condition (14) holds, then one can identify the $j$-th column of matrix $L_{i}$ as ${\rm col}_{j}(L_{i})=X_{f}\delta_{T}^{k}$. As it can be deduced from Remark 3.5 in Li et al. (2025), one can identify the whole matrix $L$ using the following formula
$L=X_{f}\odot_{\mathcal{B}}(U_{p}*X_{p})^{\top},$
where $\odot_{\mathcal{B}}$ is the Boolean product of matrices (that acts as the regular matrix product with the multiplication replaced by the AND and the addition replaced by the OR). Alternatively, by resorting to the standard matrix product, one can deduce $L$ as follows:
$L=X_{f}(U_{p}*X_{p})^{\top}\left({\rm diag}(\mathbbold{1}_{N}^{\top}X_{f}(U_{p}*X_{p})^{\top})\right)^{-1}.$
Finally, as $X_{p}$ is of full row rank (see the proof of Lemma 2), it is immediate to obtain $H$ as
$H=Y_{p}X_{p}^{\#},$
where $X_{p}^{\#}\triangleq X_{p}^{\top}(X_{p}X_{p}^{\top})^{-1}$ is a right inverse of $X_{p}$.
To make the following analysis meaningful, from now on we assume that the collected data are not informative for identifiability of the considered BCN, namely ${\mathcal{B}}_{d}\supsetneq\{(L,H)\}$, and we investigate under what conditions it is still possible to use them to evaluate certain properties or to solve some control problems. By adhering to the approach initiated in the papers by van Waarde, Eising et al. (see, e.g., van Waarde et al. (2020, 2023)) and first explored for logical networks in (Li et al., 2025, Definition 2.1), when the available data do not allow identifying the BCN that generated them, the best we can do is to understand if a certain property holds or some control problem is solvable for all the BCNs in ${\mathcal{B}}_{d}$. When so, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for the property or for the problem solution.
4 Reachability property
In this section, we provide an informativity-based characterization of the reachability of a set of states for a BCN. Before analyzing the problem in a data-driven context, we first recall the model-based approach to the reachability property.
Definition 4.
Cheng et al. (2011) Given a BCN as in (6a), we say that a state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if there exist $\tau\in\mathbb{Z}_{+}$ and an input ${\bf u}(t),t\in[0,\tau-1],$ that leads the state trajectory from ${\bf x}(0)={\bf x}_{0}$ to ${\bf x}(\tau)={\bf x}_{f}$. A set ${\mathcal{X}}\subset{\mathcal{L}}_{N}$ is reachable from ${\bf x}_{0}$, if there exists ${\bf x}_{f}\in\mathcal{X}$ that is reachable from ${\bf x}_{0}$. A state ${\bf x}_{f}$ (a set $\mathcal{X}$}) is globally reachable if it is reachable from every ${\bf x}_{0}$. The BCN (6a) is reachable if every state ${\bf x}_{f}$ is globally reachable.
The reachability of a single state, of a set $\mathcal{X}$} and of the whole BCN can be characterized (see, e.g., Section 3 in Fornasini & Valcher (2013b) and Chapter 16 in Cheng et al. (2011)) by resorting to the matrix
$L_{tot}\triangleq L_{1}\vee L_{2}\vee\dots\vee L_{M}\in{\mathcal{B}}^{N\times N},$ (15)
and the associated digraph $\mathcal{D}(L_{tot})\triangleq(\mathcal{V},\mathcal{E}_{tot})$. The state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if and only if there exists $\tau\in\mathbb{Z}_{+}$ such that $L_{tot}^{\tau}]_{ij}=1$, or equivalently, there is a path of length $\tau$ from $j$ to $i$ in $\mathcal{D}(L_{tot})$. [^3] In the sequel, we will interchangeably write that a node (in the digraph) or a state (of the BCN) is reachable from another node or state. More generally, as there is a bijective correspondence between states and nodes, with a slight abuse of terminology, we will often refer to state $\delta^{i}_{N}$ to denote node $i$ in the digraph, and vice versa. This is the case if and only if $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)\delta^{j}_{N}>0$. Consequently, the state ${\bf x}_{f}=\delta^{i}_{N}$ (the set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$) is globally reachable if and only if all entries of the row vector $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)$ (of the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)$) are positive. This implies that a BCN is reachable if and only if the matrix $\sum_{\tau=0}^{N-1}L_{tot}^{\tau}$ has all positive entries, which means that $L_{tot}$ is irreducible, or equivalently $\mathcal{D}(L_{tot})$ is strongly connected.
We are now ready to introduce the notion of informativity for reachability of a set of states $\mathcal{X}$ that extends Definition 3.7 in Li et al. (2025).
Definition 5.
Given a set ${\mathcal{X}}\subset\mathcal{L}_{N}$ (in particular, a state ${\bf x}_{f}\in{\mathcal{L}}_{N}$), we say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for reachability of $\mathcal{X}$ if the set $\mathcal{X}$ (the state ${\bf x}_{f}$) is globally reachable for all the BCNs in ${\mathcal{B}}_{d}$.
Since the reachability of some state ${\bf x}_{f}$ from another state ${\bf x}_{0}$ is related to the existence of a finite number of state transitions that lead the state from ${\bf x}_{0}$ to ${\bf x}_{f}$, it is clear that when we try to verify this property for all BCNs in ${\mathcal{B}}_{d}$, we can rely only on the state transitions that are revealed by the data. This can be performed by first identifying the best possible approximation of $L_{tot}$ based on data, say $L_{tot}^{d}\in{\mathcal{B}}^{N\times N}$, and then verifying the reachability properties as explained in the model-based approach by referring to such matrix. Note that in general (differently from $L_{tot}$) the matrix $L_{tot}^{d}$ may have zero columns. This happens if there exists $j\in[1,N]$ such that $\delta^{j}_{N}$ does not appear as a column of $X_{p}$. We have the following result.
Proposition 6.
Set $L_{tot}^{d}\triangleq X_{f}\odot_{\mathcal{B}}X_{p}^{\top}$. Then for every $(i,j)\in[1,N]\times[1,N]$, there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{i}\end{bmatrix}$, (16)
if and only if $L_{tot}^{d}]_{ij}=1$. Therefore the data are informative for reachability of the set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$ (of the state ${\bf x}_{f}=\delta^{i}_{N}$) if and only if all entries of the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$ (of the row vector $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$) are positive.
Proof.
The first statement follows from Lemma 3.8 in Li et al. (2025). The second part is immediate from the first statement and the comments before Proposition 6. ∎
Proposition 6 provides an extension of Theorem 3.9 in Li et al. (2025), where a characterization of informativity for (the overall) reachability is provided. Indeed, differently from Li et al. (2025), we focus on the reachability of a set or a state, since this will be useful in the following sections. Moreover, as an interesting byproduct, we have deduced the matrix $L_{tot}^{d}$ that represents the best possible “under-approximation" of the matrix $L_{tot}$, in the sense that it is the largest Boolean matrix that satisfies $L_{tot}^{d}\leq L_{tot}$ for all the matrices $L_{tot}$ of the BCNs in ${\mathcal{B}}_{d}$.
It is worth underlining that Proposition 6 only provides an answer to the question of whether or not data are informative for reachability, but it does not offer any insight into the way one can select an input sequence that achieves the goal. To this end, we introduce an algorithm, which also allows us to explore another concept that will be useful in the subsequent analysis.
Definition 7.
Given a BCN described as in (6a), the basin of attraction ${\mathcal{S}}({\bf x}_{f})$ of a state ${\bf x}_{f}\in{\mathcal{L}}_{N}$ is the set of states ${\bf x}_{0}\in{\mathcal{L}}_{N}$ from which it is possible to reach ${\bf x}_{f}$.
Clearly, a state ${\bf x}_{f}$ is globally reachable if and only if its basin of attraction ${\mathcal{S}}({\bf x}_{f})$ coincides with ${\mathcal{L}}_{N}$.
Algorithm 1, below, provides a way to determine the basin of attraction $\mathcal{S}$$^{*}$ of a set $\mathcal{X}$}, by this meaning the union of the basins of attractions of the states in $\mathcal{X}$}. For each state in ${\mathcal{S}}^{*}\setminus{\mathcal{X}}$ the algorithm also returns a possible choice for the input that if applied to that state at a certain time instant $\bar{t}$, it ensures that at the next time instant $\bar{t}+1$ the distance between the state and the target set $\mathcal{X}$ decreases. Therefore, it readily follows that, by applying the inputs provided by the algorithm, if the set $\mathcal{X}$ is globally reachable, it will be reached in a finite number of steps.
Algorithm 1 Reachability of $\mathcal{X}$}
Input: - The data matrices $X_{p},X_{f}$ and $U_{p}$;
     - The set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$.
Output: - A possible choice of inputs $({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})$;      - The basin of attraction ${\mathcal{S}}^{*}\triangleq\bigcup_{{\bf x}\in{\mathcal{X}}}\mathcal{S}({\bf x})$ of
      $\mathcal{X}$}.
Initialization: $({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
Set $d=0$, ${\mathcal{S}}_{0}=\mathcal{X}$.
Iterative procedure:
if ${\mathcal{S}}_{d}\neq\emptyset$, then
  • $d\leftarrow d+1$
  • ${\mathcal{S}}_{d}=\emptyset$
  • for $k\in[1,T]$, do
       if
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix},\qquad\exists\delta^{q}_{N}\in{\mathcal{S}}_{d-1},$
    and $\delta^{r}_{N}\not\in\cup_{i=1}^{d}{\mathcal{S}}_{i}$, then
    ${\mathcal{S}}_{d}\leftarrow{\mathcal{S}}_{d}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}\triangleq U_{p}\delta_{T}^{k}$
       end if
    end for
  • Go back to the Iterative procedure.
else set $\mathcal{S}^{*}=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$
end if
Conclusion: Return ($({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})$, $\mathcal{S}^{*}$)
Note that Algorithm 1 identifies first the set ${\mathcal{S}}_{1}$ of states in ${\mathcal{L}}_{N}\setminus\mathcal{X}$ that have a successor in $\mathcal{S}_{0}=\mathcal{X}$. For each such state, say $\delta^{r}_{N}$, it memorizes in ${\bf u}_{r}$ one of the values of the input that allows the transition from $\delta^{r}_{N}$ to one of the states in $\mathcal{X}$, say $\delta^{q}_{N}$ (referring to the algorithm notation). Then it identifies the set ${\mathcal{S}}_{2}$ of the states that do not belong to ${\mathcal{S}}_{0}\cup{\mathcal{S}}_{1}=\color[rgb]{0,0,0}{\mathcal{X}}\cup{\mathcal{S}}_{1}$, but have a successor in ${\mathcal{S}}_{1}$. Again, for each $\delta^{r}_{N}\in{\mathcal{S}}_{2}$, it memorizes one of the values of the input that allows the transition from $\delta^{r}_{N}$ to some state in ${\mathcal{S}}_{1}$. Since the set of states of a BCN is finite, there exists $d\geq 1$ such that ${\mathcal{S}}_{d}=\emptyset$. If at that stage $\bigcup_{i=0}^{d-1}{\mathcal{S}}_{i}=\mathcal{L}_{N}$, this means that $\mathcal{X}$ is globally reachable.
5 Equilibrium points of a BCN
We now recall the notions of equilibrium point and limit cycle in the context of logical systems.
Definition 8.
Given a BN as in (5), a state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ is an equilibrium point of the BN if ${\bf x}_{e}=L{\bf x}_{e}$. Given a BCN as in (6a), we say that ${\bf x}_{e}$ is an equilibrium point of the BCN corresponding to the input ${\bf u}=\delta^{i}_{M}$ if ${\bf x}_{e}$ is an equilibrium point of its $i$-th subsystem (8), i.e., ${\bf x}_{e}=L\ltimes\delta^{i}_{M}\ltimes{\bf x}_{e}$.
Note that ${\bf x}_{e}=\delta^{j}_{N}$ is an equilibrium point of a BN or a BCN if and only if the node $j$ has a self-loop in ${\mathcal{D}}(L)$ or ${\mathcal{D}}(L_{tot})$, respectively. This amounts to saying that $L]_{jj}$ or $L_{tot}]_{jj}$, respectively, is unitary.
Based on the comments about $L_{tot}$ in the previous section, it is immediate to deduce that the set of equilibrium points compatible with the available data, say, $\mathcal{X}_{e}^{d}$, namely the set of states that are equilibria in all the BCNs in $\mathcal{B}_{d}$, coincides with the set of states $\delta^{j}_{N}$ such that $L_{tot}^{d}]_{jj}=1$, or, equivalently, $X_{f}(X_{p})^{\top}]_{jj}>0$, which is consistent with the analysis proposed in Section 3.2 of Li et al. (2025). As an alternative, we can notice that $L_{tot}^{d}]_{jj}=1$ if and only if $\exists k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{j}\end{bmatrix}\Leftrightarrow(X_{p}\odot X_{f})\delta^{k}_{T}=\delta_{N}^{j},$
where we recall that $X_{p}\odot X_{f}$ is the Hadamard product of $X_{p}$ and $X_{f}$. This characterization has the advantage of allowing us to identify the input to which the equilibrium point corresponds as ${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}=\delta_{M}^{i},\exists i\in[1,M]$.
Since the number of possible states is finite, as time evolves, the state of a Boolean network either becomes constant (and hence reaches an equilibrium point) or it becomes periodic. In the latter case, we talk about a limit cycle. Note that an equilibrium point is actually a limit cycle of unit length.
Definition 9.
Fornasini & Valcher (2013b) Given a BN as in (5), an ordered sequence of distinct logical vectors $(\delta_{N}^{i_{1}},\delta_{N}^{i_{2}},\dots,\delta_{N}^{i_{k}})$ is a limit cycle $\mathcal{C}$ if, taken ${\bf x}(0)=\delta_{N}^{i_{\ell}}$ for some $\ell\in[1,k]$, the corresponding state evolution ${\bf x}(t)$ is periodic of period $k$ and, for every $t\in\mathbb{Z}_{+}$, ${\bf x}(t)=\delta_{N}^{i_{j}}$, with $j\in[1,k]$ satisfying $j=(t+\ell){\rm mod}k$.
We will make use of this concept in Section 7.
6 Data-driven safe control
In Fornasini & Valcher (2016) the safe control of a BCN was first addressed, then it was extended to handle random impulsive logical control networks in Zhou et al. (2023) and Probabilistic Boolean control networks using an event-triggered approach in Shao et al. (2025). The safe control problem is stated as follows. Given a BCN (6a), suppose that the set of states ${\mathcal{L}}_{N}$ can be partitioned into a set of unsafe states ${\mathcal{X}}_{u}$ and a set of safe states ${\mathcal{X}}_{s}\triangleq{\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}$. Under what conditions is it possible to design a control input so that every state trajectory that originates from a safe state remains indefinitely in ${\mathcal{X}}_{s}$ and every trajectory that starts from an unsafe state enters ${\mathcal{X}}_{s}$ in a finite number of steps and there remains?
In Fornasini & Valcher (2016) a complete characterization of the problem solvability has been provided.
Proposition 10.
Given a BCN (6a) and the set ${\mathcal{X}}_{u}$ of unsafe states, the safe control problem is solvable if and only if
  • i)
    for every ${\bf x}\in{\mathcal{X}}_{s}={\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}$ there exists ${\bf u}\in{\mathcal{L}}_{M}$ such that $L\ltimes{\bf u}\ltimes{\bf x}\in{\mathcal{X}}_{s}$;
  • ii)
    the set ${\mathcal{X}}_{s}$ is reachable from every ${\bf x}\in{\mathcal{X}}_{u}$, which amounts to saying that for every ${\bf x}\in{\mathcal{X}}_{u}$ there exists $\bar{\bf x}\in{\mathcal{X}}_{s}$ such that $\bar{\bf x}$ is reachable from $\bf x$}.
Moreover, if the safe control problem is solvable, then it is solvable by means of a state feedback law.
Remark 11.
It is worth remarking that safe control to a set ${\mathcal{X}}_{s}$ is similar to the problem of stabilizing a BCN to the set ${\mathcal{X}}_{s}$, but it does not coincide with it. Indeed, while stabilization only requires that every state trajectory eventually enters ${\mathcal{X}}_{s}$, safe control imposes the additional constraint that states belonging to ${\mathcal{X}}_{s}$ can never leave ${\mathcal{X}}_{s}$, not even in a transient phase.
The data-driven version of the safe control problem is the following one.
Problem 1.
Given the data $({\bf x}_{d}$, ${\bf u}_{d})$ and the unsafe set ${\mathcal{X}}_{u}\subsetneq{\mathcal{L}}_{N}$, determine (if it exists) a state feedback control law ${\bf u}(t)=K{\bf x}(t)$, $K\in\mathcal{L}_{M\times N}$, that solves the safe control problem for every BCN compatible with the data.
When Problem 1 is solvable, the data $({\bf x}_{d}$, ${\bf u}_{d})$ will be called informative for safe control (with respect to the unsafe set ${\mathcal{X}}_{u}$). In the following theorem we provide necessary and sufficient conditions for the solvability of Problem 1.
Theorem 12.
Consider the unsafe set ${\mathcal{X}}_{u}$ and set ${\mathcal{X}}_{s}={\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$. Problem 1 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d})$ are informative for safe control with respect to the unsafe set ${\mathcal{X}}_{u}$) if and only if the following two conditions hold:
  • i)
    for every $\delta_{N}^{j}\in{\mathcal{X}}_{s}$ there exist $k\in[1,T]$ and $\delta_{N}^{\ell}\in{\mathcal{X}}_{s}$ such that
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}$; (17)
  • ii)
    the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$ has (all) positive entries.
Proof.
Problem 1 is solvable if and only if conditions i) and ii) in Proposition 10 are verified for all the BCNs compatible with the data.
In order to check condition i) of Proposition 10, we need to identify from data if each state in ${\mathcal{X}}_{s}$ has (at least) one successor in ${\mathcal{X}}_{s}$. This amounts to verifying if for every $\delta^{j}_{N}\in{\mathcal{X}}_{s}$ there exists $k\in[1,T]$ and $\delta^{\ell}_{N}\in{\mathcal{X}}_{s}$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}$ (18)
which is exactly i) in the theorem statement. If this is the case, an input that allows the transition is
${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}.$ (19)
On the other hand, by Proposition 6, condition ii) in Proposition 10 is trivially equivalent to condition ii) in the theorem. ∎
Theorem 12 establishes data-based conditions for the solvability of Problem 1; however, it does not yield an explicit solution, which is instead provided in Algorithm 2, below. In Algorithm 2 we first check condition i) of Theorem 12. If such condition is not verified, the algorithm stops and provides a negative outcome. If, on the other hand, the first check is successful, it performs the reachability check required by condition ii) from each state in ${\mathcal{X}}_{u}$ according to the same logic as in Algorithm 1.
Algorithm 2 returns, in Step 1, an input that keeps states already in $\mathcal{X}_{s}$ within $\mathcal{X}_{s}$, and in Step 2, an input that drives states in $\mathcal{X}_{u}$ closer to $\mathcal{X}_{s}$. Such inputs are used to build a state feedback matrix $K$ that solves the problem.
Algorithm 2 Safe control
Input: - The data matrices $X_{p},X_{f}$ and $U_{p}$;
     - The set ${\mathcal{X}}_{s}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$.
Output: - Is Problem 1 solvable? Yes/No.
     - A state feedback matrix $K$ that solves the
      problem.
Initialization: $({\bf u}_{1}^{1},{\bf u}_{2}^{1},\dots,{\bf u}_{N}^{1})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$,
$({\bf u}_{1}^{2},{\bf u}_{2}^{2},\dots,{\bf u}_{N}^{2})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
  1. 1.
    • Set $\mathcal{S}_{0}=\emptyset$.
    • for $k\in[1,T]$, do
        if
      $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix}$,
      with $\delta^{r}_{N}\in{\mathcal{X}}_{s}\setminus{\mathcal{S}}_{0}$ and $\delta^{q}_{N}\in{\mathcal{X}}_{s}$, then
      ${\mathcal{S}}_{0}\leftarrow{\mathcal{S}}_{0}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}^{1}\triangleq U_{p}\delta_{T}^{k}$
         end if
      end for
      if
      $\mathcal{S}_{0}\not\equiv\mathcal{X}_{s}$, then
          return (‘No’, ‘None’).
      end if
  2. 2.
    • Apply Algorithm 1 with respect to the set $\mathcal{X}_{s}$.
      Let ($({\bf u}_{1}^{2},{\bf u}_{2}^{2},\dots,{\bf u}_{N}^{2})$, $\mathcal{S}^{*}$) be the corresponding outcome.
    • for $i\in[1,N]$, do
           ${\bf u}_{i}={\bf u}_{i}^{1}+{\bf u}_{i}^{2}$
      end for
    • Set $K=\begin{bmatrix}{\bf u}_{1}&\dots&{\bf u}_{N}\end{bmatrix}$.
  3. 3.
    if $\mathcal{S}^{*}=\mathcal{L}_{N}$, then
        return (‘Yes’, $K$)
    else,
      return (‘No’, ‘None’)
    end if
Example 13, below, explores on a small size network the results of this section pertaining data-driven safe control.
Example 13.
Consider a BCN (6a) with $N=7$ and $M=3$, described by the following matrix
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\big[\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ \delta_{7}^{5}\ \delta_{7}^{2}\ \delta_{7}^{7}\ \delta_{7}^{5}\ |\$
$|\ \delta_{7}^{6}\ \delta_{7}^{1}\ \delta_{7}^{3}\ \delta_{7}^{2}\ \delta_{7}^{4}\ \delta_{7}^{5}\ \delta_{7}^{7}\ |\ \delta_{7}^{7}\ \delta_{7}^{6}\ \delta_{7}^{2}\ \delta_{7}^{3}\ \delta_{7}^{1}\ \delta_{7}^{6}\ \delta_{7}^{6}\big].$
The equilibrium points of this BCN are $\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{6},\delta_{7}^{7}\}$. We perform a single ($r=1$) offline experiment in the time interval $[0,T]$, with $T=12$, and we collect the following data:
$X_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{7}^{1}&\delta_{7}^{7}&\delta_{7}^{7}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{4}&\delta_{7}^{3}&\delta_{7}^{3}&\delta_{7}^{2}\end{array}\right],$
$U_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}\end{array}\right],$
$X_{f}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c}\delta_{7}^{7}&\delta_{7}^{7}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{6}&\delta_{7}^{5}&\delta_{7}^{1}&\delta_{7}^{4}&\delta_{7}^{3}&\delta_{7}^{3}&\delta_{7}^{2}&\delta_{7}^{2}\end{array}\right].$
The BCNs compatible with these data can be generically represented by the following matrix $\tilde{L}$
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\big[\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ *\ *\ *\ *\ |\$
$|\ \delta_{7}^{6}\ *\ \delta_{7}^{3}\ *\ *\ \delta_{7}^{5}\ \delta_{7}^{7}\ |\ \delta_{7}^{7}\ *\ *\ \delta_{7}^{3}\ \delta_{7}^{1}\ *\ \delta_{7}^{6}\big].$
where the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{7}$). Since
$L_{tot}^{d}=\begin{bmatrix}\delta^{4}_{7}+\delta^{6}_{7}+\delta^{7}_{7}&\ \ \delta^{2}_{7}&\ \ \delta^{2}_{7}+\delta^{3}_{7}&\ \ \delta^{3}_{7}&\ \ \delta^{1}_{7}&\ \delta^{5}_{7}&\ \delta^{6}_{7}+\delta^{7}_{7}\end{bmatrix}$
we obtain that the set of the equilibrium points compatible with all the BCNs in $\mathcal{B}_{d}$ is $\mathcal{X}_{e}^{d}=\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{7}\}.$
Assume now that we want to implement a state feedback safe control with respect to the unsafe set $\mathcal{X}_{u}=\{\delta_{7}^{3},\delta_{7}^{4},\delta_{7}^{7}\}$. We need to understand if the data are informative for safe control with respect to $\mathcal{X}_{u}$, or, equivalently, if conditions i) and ii) of Theorem 12 are satisfied. To this end, we can apply Algorithm 2, where in Step 1 we can readily verify that for every state in $\mathcal{X}_{s}=\mathcal{L}_{7}\setminus\mathcal{X}_{u}=\{\delta_{7}^{1},\delta_{7}^{2},\delta_{7}^{5},\delta_{7}^{6}\}$ there exists a transition ending in a state that is still inside $\mathcal{X}_{s}$, namely condition i) of Theorem 12 holds. In Step 2, we instead examine condition ii) by employing Algorithm 1 to determine the reachability of $\mathcal{X}_{s}$ in every BCN in $\mathcal{B}_{d}$. We obtain: $\mathcal{S}_{0}=\mathcal{X}_{s}$, $\mathcal{S}_{1}=\{\delta_{7}^{3},\delta_{7}^{7}\}$, and $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, and hence $\mathcal{S}^{*}=\mathcal{L}_{7}$. This means that also condition ii) of Theorem 12 is satisfied, and thus the data are informative for safe control with respect to $\mathcal{X}_{u}$. Algorithm 2 also provides the state feedback matrix $K=\begin{bmatrix}\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}$.
7 Data-driven output regulation
We now address the second problem considered in this paper, namely output regulation, which consists in designing a control input such that the output of a BCN becomes constant and equal to a desired value, after a finite number of steps. More specifically, we pursue this goal by designing a control action in the form of a state feedback.
Before addressing the problem in a data-driven framework, we review its definition and model-based solution (see, e.g., Fornasini & Valcher (2013a, 2014)).
Definition 14.
Given a BCN (6), the output regulation problem to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback is said to be solvable if there exists $K\in\mathcal{L}_{M\times N}$ such that for every ${\bf x}(0)\in\mathcal{L}_{N}$, ${\bf y}(t)={\bf y}^{*},\forall t\geq\tau$, $\exists\tau\in\mathbb{Z}_{+}$, when ${\bf u}(t)=K{\bf x}(t).$
The first step of the model-based solution consists in finding all the states that generate the desired output $\bf y$$^{*}$, namely the states belonging to the set $\mathcal{X}({\bf y}^{*})\triangleq\{\delta_{N}^{j}:H\delta_{N}^{j}={\bf y}^{*}\}$. Once we have determined this set, we have to ensure that, under some state feedback law ${\bf u}(t)=K{\bf x}(t),\exists K\in\mathcal{L}_{M\times N}$, the state evolution of the resulting BN, described by
${\bf x}(t+1)=L_{K}{\bf x}(t),$ (25)
where $L_{K}\triangleq L\ltimes K\ltimes\Phi_{N}\in\mathcal{L}_{N\times N},$ and $\Phi_{N}$ is the power-reducing matrix Cheng et al. (2011), satisfying $\Phi_{N}{\bf x}(t)={\bf x}(t)\ltimes{\bf x}(t)$, is eventually constrained within $\mathcal{X}({\bf y}^{*})$. This imposes that all limit cycles (in particular, equilibrium points) of the BN (25) have states inside $\mathcal{X}({\bf y}^{*})$.
Since the effect of state feedback on a BCN is that of selecting, for each state, only one of its possible successors, corresponding to a specific choice of the input, each limit cycle appearing in the BN (25) corresponds to a periodic state trajectory of the BCN (6). On the other hand, the periodic state trajectories of the BCN (6) coincide with the cycles appearing in the digraph ${\mathcal{D}}(L_{tot})$. So, if we denote by $\mathcal{C}({\bf y}^{*})$ the set of all cycles in ${\mathcal{D}}(L_{tot})$ whose nodes correspond to states belonging to $\mathcal{X}({\bf y}^{*})$, we first need to verify that $\mathcal{C}({\bf y}^{*})\neq\emptyset$, and then
that it is globally reachable, meaning that the union of the basins of attraction of the states in $\mathcal{C}({\bf y}^{*})$ covers the whole state space $\mathcal{L}_{N}$.
In the following proposition, we formalize the necessary and sufficient conditions for solving the output regulation problem by state feedback in a model-based set-up.
Proposition 15.
Fornasini & Valcher (2014) Given a BCN as in (6) and a desired output value ${\bf y}^{*}\in{\mathcal{L}}_{P}$, the problem of output regulation to $\bf y$$^{*}$ by state feedback is solvable if and only if
  • i)
    $\mathcal{C}({\bf y}^{*})\neq\emptyset$;
  • ii)
    $\mathcal{C}({\bf y}^{*})$ is globally reachable, i.e., $\cup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}{\mathcal{S}}({\bf x})\!=\!{\mathcal{L}}_{N}$.
The data-driven version of the output regulation problem is given below.
Problem 2.
Given the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$, and some desired value ${\bf y}^{*}\in{\mathcal{L}}_{P}$ for the output, determine (if possible) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that, by applying ${\bf u}(t)=K{\bf x}(t)$, for every initial condition ${\bf x}(0)$, the output trajectories of all the BCNs compatible with the data satisfy ${\bf y}(t)={\bf y}^{*}$ for every $t\geq\tau,\exists\tau\in\mathbb{Z}_{+}$.
When Problem 2 is solvable, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback.
In order to solve Problem 2, we first determine from data the set of states that generate the desired output $\bf y$$^{*}$ in all BCNs in ${\mathcal{B}}_{d}$, i.e.,
$\mathcal{X}^{d}({\bf y}^{*})\!=\!\left\{\delta_{N}^{j}\in{\mathcal{L}}_{N}:\exists k\in[1,T]\ \text{s.t.}\begin{bmatrix}X_{p}\cr Y_{p}\end{bmatrix}\delta_{T}^{k}=\!\begin{bmatrix}\delta_{N}^{j}\cr{\bf y}^{*}\end{bmatrix}\right\}\!$.
To identify this set, we can proceed as in Algorithm 3.
Algorithm 3 Construction of $\mathcal{X}^{d}({\bf y}^{*})$
Input: - The data matrices $X_{p}$ and $Y_{p}$;
     - ${\bf y}^{*}\in\mathcal{L}_{P}$.
Output: The set $\mathcal{X}^{d}({\bf y}^{*})$ of all states of the BCNs in
      ${\mathcal{B}}_{d}$ that generate the desired output $\bf y$$^{*}$.
Initialization: Set $\mathcal{X}^{d}({\bf y}^{*})=\emptyset$.
  1. 1.
    Compute
    ${\bf v}^{\top}=({\bf y}^{*})^{\top}Y_{p}\in\mathcal{B}^{1\times T},$
    and note that ${\bf v}]_{k}\neq 0$ if and only if $X_{p}\delta_{T}^{k}\in\mathcal{X}^{d}({\bf y}^{*})$.
  2. 2.
    Define $\mathcal{K}\triangleq\{k_{1},k_{2},\dots,k_{y}\}=\$nonzero($\bf v$})
  3. 3.
    if $\mathcal{K}\neq\emptyset$, then
      for $k\in\mathcal{K}$, do
        if $X_{p}\delta_{T}^{k}\notin\mathcal{X}^{d}({\bf y}^{*})$, then
          $\mathcal{X}^{d}({\bf y}^{*})\leftarrow\mathcal{X}^{d}({\bf y}^{*})\cup\{X_{p}\delta_{T}^{k}\}$
        end if
      end for
    end if
  4. 4.
    Return $\mathcal{X}^{d}({\bf y}^{*})$.
At this point, we need to verify if there exist cycles whose nodes belong to $\mathcal{X}^{d}({\bf y}^{*})$. The first step consists in determining all the transitions appearing in the data involving (both initial and final) states that belong to $\mathcal{X}^{d}({\bf y}^{*})$. Based on these transitions, we then construct a digraph, whose set of nodes is $\mathcal{X}^{d}({\bf y}^{*})$, and with edges corresponding to the transitions within $\mathcal{X}^{d}({\bf y}^{*})$ identified from data. We call $L_{tot}^{d}({\bf y}^{*})$ the Boolean (but not necessarily logical) matrix associated to such a digraph. The matrix $L_{tot}^{d}({\bf y}^{*})$ represents the data-based estimate of the principal submatrix of $L_{tot}$ in (15) obtained by selecting the rows and the columns corresponding to the states in $\mathcal{X}^{d}({\bf y}^{*})$. In Algorithm 4, we exploit the well-know Johnson’s Algorithm (see Johnson (1975)) to identify the set of all cycles in ${\mathcal{D}}(L_{tot}^{d}({\bf y}^{*}))$, say $\mathcal{C}^{d}({\bf y}^{*})$.
Algorithm 4 Cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$
Input: - The data matrix $X$;
     - The set $\mathcal{X}^{d}({\bf y}^{*})=\{\delta^{i_{1}}_{N},\dots,\delta^{i_{s}}_{N}\}$.
Output: The set $\mathcal{C}^{d}({\bf y}^{*})$ of all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$.
Initialization: Set:
  • $P\triangleq\begin{bmatrix}\delta_{N}^{i_{1}}&\delta_{N}^{i_{2}}&\dots&\delta_{N}^{i_{s}}\end{bmatrix}\in{\mathcal{L}}_{N\times s}$;
  • $\tilde{\mathcal{V}}\triangleq\{i_{1},i_{2},\dots,i_{s}\}\subset[1,N]$;
  • $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$;
[…]
  1. 1.
    for $(i,j)\in[1,s]\times[1,s]$, do
    if $\exists k\in[1,T]$ s.t. $P^{\top}X_{p}]_{ik}[P^{\top}X_{f}]_{jk}\neq 0$, then
         $\tilde{\mathcal{E}}\leftarrow\tilde{\mathcal{E}}\cup\{(i,j)\}$
    end if
    end for
  2. 2.
    Apply Johnson’s Algorithm in Johnson (1975) to find all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$, say $\tilde{\mathcal{C}}_{1},\dots,\tilde{\mathcal{C}}_{n_{c}}$, $n_{c}\in\mathbb{Z}_{+}$.
  3. 3.
    Remap all the nodes of $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$ to the corresponding nodes in $[1,N]$, by using the bijective correspondence given by the matrix $P$, i.e., $1\leftrightarrow i_{1},\dots,s\leftrightarrow i_{s}$.
  4. 4.
    Define $\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}$ the remapped limit cycles.
  5. 5.
    Return $\mathcal{C}^{d}({\bf y}^{*})$.
For each edge $(i,j)$ belonging to some cycle $\mathcal{C}$$_{\ell}$ in $\mathcal{C}^{d}({\bf y}^{*})$ it is also possible to determine one of the inputs associated with such transition within the cycle, say ${\bf u}_{i}^{\ell}$. Indeed, if there exists an edge $(i,j)$, it means that the transition $i\to j$ is captured by the collected data, or, equivalently, there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{i}\cr\delta_{N}^{j}\end{bmatrix}$. (26)
Therefore, the input associated to this transition within $\mathcal{C}$$_{\ell}$ is
${\bf u}_{i}^{\ell}\triangleq U_{p}\delta_{T}^{k}.$ (27)
At this point it only remains to establish if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. This can be verified using Algorithm 1, which additionally specifies the inputs to be applied to the states outside $\mathcal{C}^{d}({\bf y}^{*})$.
We can now provide the solution to Problem 2.
Theorem 16.
Problem 2 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback) if and only if the following two conditions hold:
  • i)
    $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$;
  • ii)
    ${\mathcal{S}^{*}}\triangleq\bigcup_{{\bf x}\in\mathcal{C}^{d}({\bf y}^{*})}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
Proof.
(Necessity). Assume, by contradiction, that condition i) does not hold, i.e., $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\emptyset$, and hence, by Proposition 15, for that BCN the output regulation to $\bf y$$^{*}$ by state feedback is not possible.
Assume now that $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$, but ii) does not hold, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We can construct a BCN in ${\mathcal{B}}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\mathcal{C}^{d}({\bf y}^{*})$ and $\bigcup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}\mathcal{S}({\bf x})={\mathcal{S}^{*}}\neq\mathcal{L}_{N}$. This requires to select all the state transitions, equivalently the columns of $\tilde{L}$, that are not uniquely identified by the data (and possibly also the corresponding output). If there exists $\delta^{r}_{N}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ and $H\delta^{r}_{N}\neq{\bf y}^{*}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, by Proposition 15, for each BCN obtained in this way the output regulation to $\bf y$$^{*}$ by state feedback is not possible.
(Sufficiency). It is a direct consequence of Proposition 15 and following comments. ∎
As for Theorem 12 in Section 6, Theorem 16 only provides a method to verify, using data, whether the output regulation problem is solvable. However, it is not constructive, in the sense that it does not offer a systematic procedure for explicitly computing the solution, namely the feedback matrix $K\in\mathcal{L}_{M\times N}$. Nevertheless, the matrix $K$ can be readily obtained as
$K=\begin{bmatrix}{\bf u}_{1}&\dots&{\bf u}_{N}\end{bmatrix}$
by selecting as control input ${\bf u}_{i}$ the one defined in (27) for the states $\delta^{i}_{N}$ belonging to a cycle in $\mathcal{C}^{d}({\bf y}^{*})$, and using the input ${\bf u}_{i}$ provided by Algorithm 1 for all other states $\delta^{i}_{N}$. In cases where a state belongs to multiple cycles in $\mathcal{C}^{d}({\bf y}^{*})$, it suffices to select the control input associated with any of these cycles.
Remark 17.
In this section, we have explored how to solve the output regulation problem via state feedback based on collected data. Clearly, this strategy cannot be implemented unless offline data about the state have been collected. An alternative approach, first explored in Leifeld et al. (2018), is to perform output regulation by means of output feedback. When so, one could try to achieve this goal by collecting only input and output data. The solution proposed in Leifeld et al. (2018), and inspired by results on the stabilization of Probabilistic Boolean networks, relies on an algorithm that first computes an approximation of the output prediction matrix, based on data, and then generates a sequence of concentric annular sets, each containing the output values that lie at a specific “distance" from the desired value $\bf y$$^{*}$, with respect to the output/input/output transitions captured by the data-based output prediction matrix. Finally, using these sets, the algorithm computes, when possible, an output feedback gain.
However, output feedback strategies, which are demanding even in model-based settings, become hardly applicable in a data-driven framework, where only part of the network transitions is accessible.
Example 18.
Consider a BCN (6) with $N=6$, $M=3$, and $P=2$, described by the following matrices
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\left[\begin{array}[]{c|}\delta_{6}^{2}\ \delta_{6}^{4}\ \delta_{6}^{3}\ \delta_{6}^{3}\ \delta_{6}^{6}\ \delta_{6}^{5}\end{array}\right$.
$\left.\begin{array}[]{|c|c}\delta_{6}^{1}\ \delta_{6}^{5}\ \delta_{6}^{2}\ \delta_{6}^{2}\ \delta_{6}^{6}\ \delta_{6}^{1}&\delta_{6}^{5}\ \delta_{6}^{1}\ \delta_{6}^{4}\ \delta_{6}^{5}\ \delta_{6}^{4}\ \delta_{6}^{6}\end{array}\right],$
$H$ $=$ $\left[\begin{array}[]{c}\delta_{2}^{1}\ \delta_{2}^{2}\ \delta_{2}^{2}\ \delta_{2}^{2}\ \delta_{2}^{1}\ \delta_{2}^{1}\end{array}\right].$
Assume that we want to regulate the output of the network to the value ${\bf y}^{*}=\delta_{2}^{2}$. We perform a single ($r=1$) offline experiment in the time interval $[0,9]$, and collect the data:
$X$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c|c}\delta_{6}^{6}&\delta_{6}^{6}&\delta_{6}^{1}&\delta_{6}^{2}&\delta_{6}^{5}&\delta_{6}^{4}&\delta_{6}^{2}&\delta_{6}^{4}&\delta_{6}^{3}&\delta_{6}^{3}\end{array}\right],$
$U_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c}\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}\end{array}\right],$
$Y_{p}$ $=$ $\left[\begin{array}[]{c|c|c|c|c|c|c|c|c}\delta_{2}^{1}&\delta_{2}^{1}&\delta_{2}^{1}&\delta_{2}^{2}&\delta_{2}^{1}&\delta_{2}^{2}&\delta_{2}^{2}&\delta_{2}^{2}&\delta_{2}^{2}\end{array}\right].$
We deduce $\mathcal{X}^{d}({\bf y}^{*})=\{\delta_{6}^{2},\delta_{6}^{3},\delta_{6}^{4}\}.$ The BCNs compatible with these data can be represented by the following matrix
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\left[\begin{array}[]{c|}\delta_{6}^{2}\ \delta_{6}^{4}\ \delta_{6}^{3}\ \delta_{6}^{3}\ *\ *\end{array}\right$.
$\left.\begin{array}[]{|c|c}*\ \delta_{6}^{5}\ *\ \delta_{6}^{2}\ *\ \delta_{6}^{1}&*\ *\ *\ *\ \delta_{6}^{4}\ \delta_{6}^{6}\end{array}\right],$
where, again, the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{6}$). As $X_{p}$ is of full row rank, the matrix $H$ can be uniquely recovered. Hence, by applying Algorithm 4, we obtain that the set of all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$ is $\mathcal{C}^{d}({\bf y}^{*})=\{3,2\leftrightarrow 4\},$ where $2\leftrightarrow 4$ is the cycle of length $2$ involving $\delta_{6}^{2}$ and $\delta_{6}^{4}$. We now check if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. By following the procedure outlined in Algorithm 1, we obtain $\mathcal{S}^{*}={\mathcal{L}}_{6}$. Thus, by Theorem 16, the data are informative for output regulation to $\bf y$$^{*}$ by state feedback. Building on the argument presented after Theorem 16, we obtain the following feedback matrix
$K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}\end{bmatrix}$.
8 Conclusion
In this paper we investigated how the recently introduced data informativity approach van Waarde et al. (2020) can be adapted and extended to solve fundamental control problems in Boolean control networks. More specifically, we focused on safe control and output regulation problems, and we provided necessary and sufficient conditions, based only on previously collected data (supposed not to be informative for identification), for their solvability. Moreover, we also obtained from data the feedback matrices expressions. We validated our results by means of examples. It is worth noting that, due to the inherent complexity of these systems, BCN problems are NP-hard with respect to the state, input and output dimensions $n$, $m$ and $p$, or equivalently polynomial in the corresponding quantities $N=2^{n}$, $M=2^{m}$ and $P=2^{p}$. Importantly, the computational complexity is independent of the chosen representation: in the case of interest, the algebraic one. Moreover, relying on data instead of an exact model does not substantially change the overall computational complexity. The only difference is that it also depends on the collected data, and in particular on the number of collected data $T$, which is, however, lower-bounded by $N$. On the other hand, since data provide only partial knowledge of the system dynamics, the number of identified transitions is smaller than those appearing in the real model, and consequently, if their complete enumeration is required, it imposes a lower computational burden. Nonetheless, depending on the problem at hand, the algorithms can be easily adapted to terminate once a single solution is found, without the need to search for all feasible solutions.